### v1

# Unifying Demonstration Selection and Compression for In-Context Learning

Abstract
In-context learning (ICL) facilitates large language models (LLMs) exhibiting spectacular emergent capabilities in various scenarios. Unfortunately, introducing demonstrations easily makes the prompt length explode, bringing a significant burden to hardware. In addition, random demonstrations usually achieve limited improvements in ICL, necessitating demonstration selection among accessible candidates. Previous studies introduce extra modules to perform demonstration compression or selection independently. In this paper, we propose an ICL framework UniICL, which Unifies demonstration selection and compression, and final response generation via a single frozen LLM. Specifically, UniICL first projects actual demonstrations and inference text inputs into short virtual tokens, respectively. Then, virtual tokens are applied to select suitable demonstrations by measuring semantic similarity within latent space among candidate demonstrations and inference input. Finally, inference text inputs together with selected virtual demonstrations are fed into the same frozen LLM for response generation. Notably, UniICL is a parameter-efficient framework that only contains 17M trainable parameters originating from the projection layer. We conduct experiments and analysis over in- and out-domain datasets of both generative and understanding tasks, encompassing ICL scenarios with plentiful and limited demonstration candidates. Results show that UniICL effectively unifies 12×12$\times$12 × compression, demonstration selection, and response generation, efficiently scaling up the baseline from 4-shot to 64-shot ICL in IMDb with 24 GB CUDA allocation [^1] The code and model will be released in the final version..
Unifying Demonstration Selection and Compression for In-Context Learning
Jun Gao School of Computer Science and Technology Soochow University junegao1106@gmail.com
Figure 1: The efficiency comparison between UniICL and other compression methods in IMDb with the number of shots increasing from 0 to 64. * and the break of line chart indicates occurring memory explodes.
Figure 2: Compressing candidate demonstrations into spans of compressed virtual tokens and selecting one for the inference inputs. Compressed virtual tokens substitute the original demonstration together with the inference input, fed to the target LLM for prediction generation.
1 Introduction
In-context learning (ICL) Brown et al. (2020); Xie et al. (2021); Wang et al. (2023b) exhibits powerful performance in practical applications with the emergence of scaled-up Transformer-based Vaswani et al. (2017) Large Language Models (LLMs) Wang et al. (2023c); Yang et al. (2023); Wei et al. (2023); Wang et al. (2023a); Min et al. (2022). In ICL, the provided demonstrations activate the pre-training knowledge of LLMs to allow them to perform well on various downstream tasks such as text summarization Wang et al. (2023c); Yang et al. (2023) and text classification Min et al. (2022) without gradient updating of billion parameters. Despite its significant role in the era of LLMs, ICL also brings an enormous challenge to the input window. Specifically, inevitably introducing demonstrations directly causes length disaster Wang et al. (2024), which is more serious in long generative tasks such as question answering Ghosal et al. (2022) and text summarization Narayan et al. (2018), bringing significant memory costs and decreasing inference throughput as described in Figure 1. Except for length disaster, Liu et al. (2021) points out that the quality of selected demonstrations significantly influences the ICL performance. The mentioned two issues derive two research lines focused on selecting salient demonstrations and utilizing demonstrations efficiently.
Generally, researchers attempted to alleviate length explode Wingate et al. (2022); Mu et al. (2023); Ge et al. (2023) via soft prompts. The main idea was to train an independent compressor to compress the input into compact virtual tokens. Similarly, Liu et al. (2021); Lu et al. (2021) scores each candidate demonstration for the given inference input via an extra ranker, and Ram et al. (2023); Wang et al. (2024) finds that a fine-tuned LLM is up to the task of ranking. However, the external compressor/ranker naturally leads to the Out-of-Distribution (OoD) problem between the compressor/ranker and the target LLM Jiang et al. (2023), requiring extra large-scale data to perform alignment. In addition, the standalone compressor or ranker incurs additional memory costs as it necessitates simultaneous loading alongside the target LLM.
Motivated by the inherent semantic understanding ability of LLMs, we propose an ICL framework UniICL that unifies demonstration compression, demonstration selection, and response generation via the same frozen LLM. In UniICL, the same frozen LLM is applied thrice: it acts as the compressor to compress demonstrations into virtual tokens, as the selector to select suitable demonstrations based on the saliency measured by virtual tokens, and as the generator to produce final responses under the guidance of selected virtual demonstrations. Specifically, similar to the mask token in pretrained language understanding models Devlin et al. (2018); Liu et al. (2019), UniICL introduces the compression slot [M] and attaches $k\times[\text{M}]$ behind demonstrations to be condensed. The LLM forwards each processed demonstration and outputs corresponding $k$ hidden states on top of compression slots. Considering the latter compression slots naturally summarize the preceding demonstrations due to the transformer blocks, we train a simple projection layer to convert continuations to condensed virtual tokens. We briefly explain UniICL on the linguistic acceptability task concerning how it selects one demonstration among two compressed candidates and produces the prediction in Figure 2. On the one hand, virtual tokens effectively absorb semantic information from actual demonstrations, along with inference inputs, fed into the target LLM to perform generation as normal ICL. On the other hand, virtual tokens are again applied to measure saliency scores between demonstrations and given inference inputs within latent space, while demonstrations with higher scores are preferentially fed to the generator for response generation. Furthermore, UniICL compresses each demonstration independently of others, allowing caching virtual token sequences to configure a Demonstrations Bank (DB) for reusing, which will significantly bolster the inference throughput as demonstrated in Figure 1.
Generally, since the compressor, selector, and generator are the same frozen LLM, UniICL merely equips 17M trainable parameters originating from a projection layer and one special embedding [M], and we collect 30k public data for training. We extensively evaluate UniICL on in- and out-domain datasets of both generative and understanding tasks, covering the mainstream high- and low-resource ICL [^2] High- and low-resource ICL are distinguished by the magnitude of candidates. The former usually selects demonstrations from plentiful candidates, and the latter merely selects from a handful of data.. Results indicate that UniICL effectively substitutes 12$\times$ longer demonstrations with virtual tokens and scale up the inner LLM from 4-shot to 64-shot, significantly alleviating length disaster at an expense of 8% extra inference latency (without caching). Meanwhile, compared with the popular dense retriever, UniICL selects more suitable demonstrations in different tasks.
Our main contributions are as follows:
  • To our knowledge, we are the first to propose an ICL framework that unifies compression, selection, and generation via a single frozen LLM, inherently bypassing the OoD problem.
  • UniICL is a memory-friend framework that enables LLMs to perform large-shot ICL on consuming GPUs.
  • Simple application of virtual tokens outperforms the popular dense retriever in selecting suitable demonstrations after jointly optimized by compression and selection augmentation loss.
2 Related Work
In this section, we simply rethink the methods focus on demonstration (prompt) compression and document retrieval.
2.1 Soft Prompt Compression
Recently, researchers attempted to utilize soft prompts to convert actual tokens to dense-information virtual tokens. Mostly from a distillation perspective, Wingate et al. (2022) aligned the teacher model and the student model, where the teacher model accepted the actual task instruction while the student model fed the soft prompt. The main drawback of this approach was the lack of generalization that necessitated training for each lexically different instruction. To tackle the generalization problem, Mu et al. (2023) proposed to learn a Llama-7b to compress instruction to virtual tokens, but only compress instruction was not powerful enough since the demonstrations were much longer in practical. To compress the demonstrations, Chevalier et al. (2023) proposed AutoCompressor to generate compressed virtual tokens based on a fine-tuned LLM Zhang et al. (2022). They first randomized segmented the texts with thousands of words into model-accepted range and then recursively generated soft prompts for each segment. Similarly, Ge et al. (2023) proposed ICAE that employed a LoRA-adopted Llama-7b Touvron et al. (2023) to compress the processed demonstrations to compact virtual tokens, and then fed to a frozen LLM to perform generation. Unlike previous works, the underlying LLM in UniICL was kept frozen during training and was trained on only 30k data.
2.2 Extractive Compression
Apart from employing soft prompts, researchers also endeavored to shorten prompts by extracting informative tokens from the original ones Li (2023); Jiang et al. (2023), namely token pruning Kim et al. (2022) or token merging Bolya et al. (2022). Recent works like LLMLingua Jiang et al. (2023) and Selective Context Li (2023) shared similarities but diverged on whether to eliminate tokens with high or low Perplexity (PPL). LLMLingua emphasized tokens with high PPL, attributing them as more influential, resulting in achieving outstanding performance. As mentioned in their paper, extractive compression methods encountered Out-of-Distribution (OoD) issues between the extractor and the target LLM. To reconcile this, they fine-tuned Alpaca-7b Taori et al. (2023) using the Alpaca dataset Taori et al. (2023) to perform the alignment. However, UniICL naturally bypassed the distribution-aligned module, since the compressor and the target LLM were the same.
3 Methodology
We propose UniICL, a parameter-efficient ICL framework that unifies demonstration compression, demonstration selection, and response generation via a single LLM. Specifically, UniICL applies a general frozen LLM thrice, as a compressor to compress demonstrations into short virtual tokens, as a selector to choose suitable demonstrations for ICL, and as a generator to produce responses. As for the selection of the underlying LLM, previous work has proposed that the Decoder-only model performs better than the Encoder-Decoder model in prompt compression Mu et al. (2023). We follow this conclusion and employ popular Vicuna-7b Zheng et al. (2023) and BlueLM-7B [^3] We mainly utilize Vicuna-7B to verify the effectiveness, the experiments of BlueLM will be exhibited in Appendiex B for concise. Team (2023) as the de facto backbone in UniICL.
We present a comparison between UniICL and other recent compression methods in Table 1. Without regard to the Lora-adopted compressor, ICAE Ge et al. (2023) is most similar to UniICL in architecture and compression mode that parameter-efficiently compresses actual demonstrations into dense soft prompts.
Figure 3: Demonstration compression.
Methods Additional Compressor Compression Tool # Trainable Parameters Train Size LLMLingua YES Pruning 7B 57k AutoCompressor NO Soft Prompt 7B UNKNOWN ICAE YES Soft Prompt 70M 240k UniICL NO Soft Prompt 17M 30k
Table 1: Comparsion among recent compression methods and UniICL. Compression Tool represents the involved compression technique of different methods.
3.1 Preliminary
Empirically, a prompt may encompass in-context demonstrations and an inference input in the ICL scenario. To avoid misleading and follow the existing definitions, we assume a prompt $P=(x^{D}_{1},x^{D}_{2},...,x^{D}_{n},x^{I})$, where $x^{D}_{i}$ is the $i$th demonstrations consist of an input-output pair (e.g. [DOC] -> [Sum]), and $x^{I}$ is the inference input (e.g. documents in the summarization task). We first initialize the compression slots [M] $\in\text{R}^{d}$ from the rarely used embedding of corresponding LLM, with gradient updating during training, $d$ is the dimension of the input embeddings. The compression ratio $r$ is a hyper-parameter, which is dynamically sampled from 2 to 16 during training, to calculate the specific number of compressed slots by $k=|x^{D}|/{r}$.
3.2 Demonstration Compression
For each demonstration in $P$, UniICL activates the compression slots to aggregate information from demonstrations in the forward propagation of frozen Vicuna respectively, as illustrated in Figure 3. Practically, we first attach $k$ compression slots $=k\times\text{[M]}$M] to each demonstration $x^{D}_{i}$, formatting modified prompt fed to the Vicuna. Then, frozen Vicuna forwards the modified prompts and outputs the last hidden states $H_{i}=(h_{1},h_{2},...,h_{k})$ corresponding to $k$ compression slot, dropping the others444$\oplus$ means token-level concatenation.:
_$\_,H_{i}=\text{Encode}(x^{D}_{i}\oplus M)$ (1)
Due to the transformer block in Vicuna, $H_{i}$ is compelled to attend to the preceding actual tokens and UniICL inserts a linear layer to project $H_{i}$ into compressed virtual tokens $C^{i}=(c_{1},c_{2},...,c_{k})$:
$c_{j}=W_{p}\cdot h_{j},$ (2)
where $W_{p}$ is the parameters of the projection layer.
Figure 4: Demonstrations selection.
3.3 Demonstration Selection
Except for substituting the origin demonstrations for generation, activated virtual tokens $C^{i}$ are again applied for demonstration selection among candidates, as illustrated in Figure 4. Specifically, given an inference input $x^{I}$ and its candidate demonstrations $(x^{D}_{1},x^{D}_{2},...,x^{D}_{n})$, UniICL obtains their latent representation after average pooling:
$\bar{C^{i}}=\frac{1}{k}\sum_{j=1}^{k}c_{j},$ (3)
denote as $q$ for the input and $(k_{1},k_{2},...,k_{n})$ for candidate demonstrations for simlifying. We define the cosine similarity between $q$ and $k_{i}$ is the $i$th demonstration saliency score $S_{i}$:
$S_{i}=cos(q,k_{i}).$ (4)
3.4 In-context Generation
We employ the frozen Vicuna again to generate responses with the guiding of concatenated virtual demonstrations, as illustrated in Figure 5. For $m$-shot in-context learning, we obtain $m$ virtual demonstrations after previously mentioned slot activating and demonstration selecting, denoted as $C^{1}$ to $C^{m}$. Then, we horizontally concatenate them, keeping their relative position unmodified. Finally, the concatenated virtual demonstrations together with actual inference inputs are fed into Vicuna, performing auto-regressive generation as normal:
$y_{i}=\text{Decode}(C^{1},...,C^{m};x^{I};y_{<i})$ (5)
Except for the generative manner, virtual tokens are conveniently transferred to close-ended evaluation without modifications for understanding tasks through providing the candidate answers and measuring their perplexity (PPL)  [^5] https://huggingface.co/docs/transformers/perplexity respectively, e.g. $(ppl^{+},ppl^{-}$) for sentiment classification:
$\text{Prediction}=argmin(ppl^{+},ppl^{-}),$ (6)
where answers with PPL closest to 1 are judged to be the current prediction.
Figure 5: In-context generation.
3.5 Training
The trainable parameters in UniICL are merely 17M originating from the projection layer $W_{p}$ and the introduced compression slot [M]. We first use the common language modeling loss to optimize UniICL and get a base compression model. Then we add InfoNCE He et al. (2020) to augment the demonstration selection ability:
$\mathcal{L}=\mathcal{L}_{lm}+\mathcal{L}_{ctr}.$ (7)
Specifically, during the first training phase, we slice the source input of each training instance into two parts and randomly compress one, denoted the compression part as $x_{c}$ and the unmodified as $x_{u}$, respectively. Afterward, we attach $M$ to $x_{c}$ and get virtual tokens $C$ on top of the compression slot sequence, as described in Equ. 1 and Equ. 2. Therefore, the language modeling loss $\mathcal{L}_{lm}$ if obtained as:
$\mathcal{L}_{lm}=-\frac{1}{|y|}\sum_{t=0}logP(y_{t}|x_{2};C;y_{<t}),$ (8)
where $y=(y_{1},y_{2},...)$ is the reference label of the current training instasnce. Notably, distinguished from ICAE which always puts soft prompts in the front, UniICL keeps the relative position of $C$ and $x_{u}$ unchanged to make compressed tokens insensitive to their position in prompts. Namely, compressed tokens will be placed in the former (latter) if $x_{c}$ is in front (behind) of $x_{u}$ in the original sequence. Additionally, aiming to concatenate different virtual token sequences from different demonstrations for practical employment of ICL, we introduce concatenation compression that evenly slices the extra-long compression parts into two sub-segments during training, compresses them separately, and concatenates their virtual tokens horizontally.
Consequently, we utilize contrastive learning for selection augmentation and mine positives $x^{D}_{+}$ and negatives $x^{D}_{-}$ as illustrated in Figure 6. Specifically, given each training instance $x^{I}$ and n candidate demonstrations $(x^{D}_{1},x^{D}_{2},...,x^{D}_{n})$ from two non-crossing training subsets, we employ Vicuna to calculate the PPL concerning the golden label of $x^{I}$, denoted as $ppl^{I}$. Then, we provide the $i$th demonstration and calculate its PPL, denoted as $(ppl^{D}_{i},i\in[1,n])$. We count $ppl^{I}$ as the baseline and calculate candidate relative PPL gains:
$\widetilde{ppl^{D}_{i}}=ppl^{I}-ppl^{D}_{i},i\in[1,n].$ (9)
Hence, $k^{+}$ $(k^{-})$ is the latent representation of demonstrations that furthest reduces (increases) $ppl^{I}$ as processed in Equation 3, and the contrastive loss $\mathcal{L}_{ctr}$ can be formulated as:
$\mathcal{L}_{ctr}=\frac{exp(cos(q,k^{+}))}{exp(cos(q,k^{+}))+exp(cos(q,k^{-}))}.$ (10)
In particular, if all relative PPL gains are less than 0, namely none of the candidate demonstrations help guide Vicuna to generate the golden label, we will apply the other set of candidates.
Notably, the inner compression ratio $r$ for each training step is dynamic and randomly sampled from [2, 16] during training. After the two training phases, we blend 2,000 instances from the in-domain validation sets to evaluate the quality of virtual demonstrations provided to LLMs under various ratios. Ultimately, the preferred scheme involves compressing the demonstrations with $r=12$, more details refer to Figure 7.
Figure 6: Negatives mining pipeline.
4 Experiment
In this section, we comprehensively evaluate UniICL over out-domain datasets to verify its effectiveness. Additionally, we detailedly introduce of involved datasets and evaluation metrics in Appendiex C.
4.1 Baselines
Naive Vicuna-7b serves as the fundamental baseline fed into actual demonstrations. AutoCompressor recurrently compresses demonstrations into virtual tokens. We employ their Llama2-7b version [^6] https://github.com/princeton-nlp/AutoCompressors.. LLMLingua is a coarse-to-fine demonstration pruning method based on dropping uninformative words. We employ their released 7b version [^7] https://github.com/microsoft/LLMLingua., of which compressor is a fine-tuned Llama-2. For a meaningful comparison, we replace target LLMs of LLMLingua (GPT-3.5-Turbo or Claude-v1.3) with the Vicuna-7b. ICAE [^8] https://github.com/getao/icae. compresses demonstrations into soft prompts via a LoRA-adapted Llama2-7b, which is most similar to UniICL in model architecture and compression mode. Additionally, since selection augmentation is involved in the training of UniICL, we utilize the popular Sentence-BERT (S-BERT) Reimers and Gurevych (2019) as the dense retriever to construct an ICL pipeline for the above methods, serving as simple but effective selection-based baselines.
4.2 Settings
Considering the involved datasets and computation efficiency, we set the max allowed input length limit to 1,024 for both compression and generation. For a fair comparison, we set the allowed window of baselines to 1,024 and their compression ratio to 12. We fix the learning rate to 8e-5 and use Adam as the optimizer, and the effective batch size is 32 (8 GPUs data parallelism and 4 steps gradient accumulation). Additionally, we conducted all experiments on 8*NVIDIA A5000 24G GPUs based on BFloat 16 data type, and we set the evaluated shot to 8 for understanding tasks and 5 for generative tasks for illustration. Except for the MS MARCO preprocessed by Liang Wang et al. (2022)  [^9] https://github.com/microsoft/unilm/tree/master/simlm., the other evaluation datasets are publicly achieved at hugging face.
We apply S-BERT to pre-rank and output the top 10 similar candidates from training sets according to each inference input and employ UniICL to perform selection among them in practice due to computation efficiency for high-resource ICL. On the contrary, the low-resource ICL setting fixes the candidate demonstrations to 20 for all inference inputs, performing pre-ranking and selecting as well.
4.3 Results
Model #-shots CoLA-dev SST-2-dev IMDb ARXIV XSum Acc. R-1 R-2 R-L R-1 R-2 R-L Vicuna 0-shot 56.2 91.7 92.6 34.3 9.1 27.4 19.9 5.0 13.4 1-shot 58.2 (57.4) 90.7 (90.8) 91.9 (91.0) 34.4 (33.2) 9.1 (8.5) 27.5 (26.7) 21.2 (20.4) 5.8 (5.2) 14.5 (13.9) 2-shot 62.1 (59.8) 92.1 (91.3) 91.7 (91.7) - - - - - - 5-shot 62.3 (61.9) 93.0 (91.9) 94.1 (92.5) - - - - - - AutoCompressor 1-shot 42.1 (40.9) 85.7 (84.2) 95.0 (95.1) 27.0 (26.4) 8.4 (8.2) 26.1 (25.8) 21.3 (20.3) 6.5 (6.3) 13.7 (13.7) 2-shot 58.8 (56.3) 88.0 (86.4) 95.0 (94.6) 27.1 (26.2) 8.6 (7.9) 26.4 (25.4) 21.9 (21.4) 6.6 (6.4) 14.5 (14.1) 5-shot 59.1 (58.8) 91.3 (89.1) 94.7 (94.8) 34.5 (33.7) 9.4 (9.1) 28.7 (27.9) 22.4 (21.7) 6.9 (6.7) 14.8 (14.3) LLMLingua 1-shot 55.5 (55.0) 89.7 (89.6) 91.0 (89.9) 33.3 (33.1) 8.9 (8.7) 27.4 (27.1) 20.5 (19.7) 5.4 (5.2) 14.5 (14.4) 2-shot 56.7 (55.7) 90.7 (90.2) 91.3 (91.0) 32.9 (32.0) 8.2 (8.1) 26.9 (25.9) 20.3 (20.0) 5.2 (5.1) 14.3 (14.1) 5-shot 57.2 (56.9) 90.6 (90.2) 90.9 (91.2) 30.1 (29.7) 7.9 (7.4) 25.3 (24.6) 19.7 (18.6) 4.9 (4.9) 14.1 (14.3) ICAE 1-shot 30.9 (30.9) 61.0 (60.1) 85.7 (83.3) 26.8 (24.6) 8.2 (7.1) 24.7 (22.9) 23.5 (21.9) 8.5 (7.8) 20.9 (20.3) 2-shot 30.9 (30.9) 49.0 (52.8) 85.9 (85.9) 27.2 (25.5) 8.4 (7.6) 25.9 (24.3) 24.4 (23.2) 8.9 (8.4) 21.3 (20.8) 5-shot 30.9 (30.9) 54.2 (51.0) 85.7 (85.9) 28.3 (26.9) 8.7 (7.7) 26.6 (25.8) 25.3 (24.9) 9.2 (8.8) 22.5 (21.6) UniICL 1-shot 58.7 (58.0) 92.9 (91.7) 94.3 (92.3) 35.5 (34.7) 10.5 (10.2) 28.7 (27.9) 27.7 (25.5) 10.2 (9.1) 21.2 (20.0) 2-shot 62.4 (61.0) 92.4 (91.6) 94.9 (93.3) 36.1 (35.2) 10.8 (10.4) 29.4 (28.2) 29.4 (26.8) 11.0 (9.8) 22.3 (20.9) 5-shot 62.6 (61.8) 93.1 (92.3) 94.5 (94.0) 35.8 (35.4) 10.6 (10.2) 29.5 (28.1) 30.7 (27.6) 11.3 (10.1) 22.8 (21.4) UniICL 1-shot 59.1 (58.7) 93.0 (91.9) 94.5 (91.6) 34.8 (34.7) 10.4 (10.3) 28.1 (27.8) 29.1 (26.2) 10.8 (9.4) 22.2 (20.7) 2-shot 62.6 (61.2) 94.0 (93.0) 94.9 (92.3) 34.6 (34.3) 10.6 (10.4) 28.5 (28.3) 30.3 (28.9) 11.3 (10.5) 22.9 (21.7) 5-shot 63.3 (61.5) 94.7 (92.8) 95.0 (93.8) 35.6 (35.3) 11.0 (10.8) 29.1 (27.7) 31.1 (30.0) 11.7 (11.2) 23.5 (22.3) 8-shot 63.8 (62.6) 94.7 (93.1) 95.0 (94.2) - - - - - - UniICL + $L_{ctr}$ 1-shot 59.3 (58.9) 93.2 (92.4) 95.1 (92.8) 35.6 (35.1) 10.7 (10.5) 28.9 (28.3) 30.0 (27.9) 11.3 (10.1) 22.8 (21.5) 2-shot 62.4 (62.0) 94.5 (92.8) 94.8 (93.4) 36.8 (35.3) 10.8 (10.6) 29.6 (28.9) 30.8 (29.2) 11.4 (10.7) 23.0 (21.9) 5-shot 64.3 (61.8) 94.7 (93.4) 96.1 (94.2) 37.1 (34.9) 11.3 (11.2) 30.0 (29.3) 32.5 (30.6) 12.3 (11.8) 24.7 (23.8) 8-shot 64.7 (63.3) 94.7 (94.1) 95.6 (95.0) - - - - - -
Table 2: The high- and low-ICL results on CoLA-dev, SST-2-dev, and IMDb. Results in () represent low-resource ICL. represents the demonstrations selected by UniICL, and the others are selected by S-BERT. +$L_{ctr}$ indicates the selection augmented UniICL (optimized with Equation 7). Bold (underline) represents the best performance on high- and low-resource ICL.
We comprehensively evaluate the ICL performance of UniICL on the out-domain dataset CoLA, SST-2, and IMDb by close-ended evaluation and ARXIV by open-ended evaluation, as demonstrated in Table 2. Specifically, UniICL outperforms naive Vicuna-7b fed with actual candidate demonstrations, which indicates that virtual demonstrations are more efficient and informative for guiding the target LLM, and UniICL outperforms all the baselines by compressing the same demonstrations pre-ranked by S-BERT. Additionally, UniICL achieves further performance gains after selecting demonstrations via itself. Additionally, open-ended results indicate that virtual demonstrations still efficiently capture semantic information for ICL guiding, even though summarization demonstrations are much longer than understanding ones. Regarding ARXIV, the original ICL is not helpful enough due to its extremely over-length document, leaving little room for demonstrations. UniICL works as expected by compressing demonstrations and concatenating virtual demonstrations, and achieves +2.8 R-1 gains in the 5-shot setting that selects demonstrations via selection-augmented virtual tokens.
Furthermore, The ablation experiments of +$\mathcal{L}_{ctr}$ show that UniICL is faced with performance degradation without $L_{ctr}$ and the performance gap becomes larger with the number of demonstrations increasing. The results of BlueLM are exhibited in Appendiex B.
Passage Ranking
Method MS MARCO
BM25 18.5
Vicuna 28.9
AutoCompressor 29.3
LLMLingua 27.8
ICAE 30.2
UniICL 31.6
Table 3: Results on MS MARCO. Vicuna applies the last hidden states of [EOS] to represent sentences in latent space. Following the previous study, we report MRR@10. $\dagger$ means citing from Liang Wang et al. (2022).
Since the virtual tokens naturally summarize semantic information of preceding sequences, we evaluate UniICL on the out-domain MS MARCO dataset in Table 3. UniICL significantly outperforms the sparse retrieval method BM25 algorithm and other compression methods. Notably, we don’t compare UniICL with the popular retrieval models Reimers and Gurevych (2019); Wang et al. (2024) since most of them are fine-tuned on this dataset, unfair for comparison.
Figure 7: The overall sensitivity analysis of compression ratio.
5 Analysis
5.1 Compression Ratio
During training, the compression ratio is dynamically sampled from 2 to 16. We mix up 2,000 instances from the in-domain validation set, 1,000 for XSUM, and 1,000 for CICERO to select the compression ratio for UniICL in Figure 7, with the backbone of Vicuna and BlueLM respectively. Specifically, UniICL compresses the latter cut-off part while keeping the former ones uncompressed. Therefore, we can measure the dense information quality of the same content with different compression ratios by ROUGE-1 since it is more sensitive to token-level differences. The performance is relative smoothing when the compression ratio changes from 4×4$\times$4 × to 12×12$\times$12 ×. However, when it comes to 16×16$\times$16 ×, an obvious drop occurs. Therefore, we set the compression ratio to 12 by default and apply this ratio to all experiments. The 1024×1024$\times$1024 × compression ratio is equal to compressing anything to a single virtual token, due to the maximum allowed input length for compression being 1,024.
5.2 Efficiency Analysis
In UniICL, we incorporate an additional 17M trainable parameters into the 7b backbone, accounting for an approximate increase of 2.4%. We additionally evaluate the memory costs inference latency of UniICL and other compression methods in Figure 1. Notably, compressing each demonstration independently and allowing horizon concatenation among different virtual demonstrations convert the extra compression latency be a one-shot deal that the virtual tokens can be stored, formulating a Demonstration Bank (DB) for further reusing. In this case, UniICL will eliminate the extra latency if the selected demonstrations have been compressed and cached (UniICL+Caching). Despite this, parallel computation facilitates the compressing process, resulting in minimal throughput degradation (UniICL and Baseline). What’s more, the naive 7B LLM occurs memory explosion for 8-shot settings, while UniICL successfully scales up to 64-shots within 24GB CUDA allocation.
5.3 Necessarity of UniICL
Intuitively, UniICL applies a “heavy” LLM as the compressor, but the generator and the compressor share all parameters without loading double parameters as LLMLingua, bypassing the OoD problem between the compressor and generator. Additionally, ICAE switches the LoRA parameters for compression and generations respectively, which appear to make the compressor and the generator share parameters. However, the compressor and the generator in ICAE are still two different LLMs inherently after assembling LoRA weight. Furthermore, the continual switch also breaks the latency-free merit of LoRA, resulting in more inference latency than the projection layer in UniICL, as depicted in Figure 1.
6 Conclusion
This paper proposes UniICL, a parameter-efficient ICL framework that unifies demonstration selection, demonstration compression, and final response generation via a frozen LLM. Experimental results show that the generated virtual tokens substitute the 12$\times$ longer actual demonstrations with minimal time expenditure, scaling up the number of demonstrations from 4 to 64.
7 Limitations
Our study, while proposing an efficient unified ICL framework for demonstration compression and selection, still has limitations. Firstly, UniICL is limited to the realm of naive ICL leaving other advanced LLM prompting methods, e.g. Retrieval Augment Generation (RAG) and Chain-of-Thought (CoT) unexplored. Limited to the hardware, we employ the underlying LLM at a scale of 7 billion parameters. Larger-scale LLMs are welcome to enrich our findings in future studies.
Appendix A In-Domain Evaluation
Method XSUM CICERO R-1 R-2 R-L R-1 R-2 R-L Vicuna 19.9 5.0 13.5 17.3 3.3 14.3    +UniICL 27.2 9.2 22.5 24.6 5.1 21.7
Table 4: The in-domain results of XSUM and CICERO. Vicuna generates responses conditioned on the truncated inference input and UniICL first compresses the over-length input into virtual tokens.
We conduct the zero-shot in-domain generation evaluation on the entire test set of XSUM and CICERO in Table 4, by compressing the latter half to virtual tokens and keeping the former unmodified. UniICL significantly outperforms the baseline Vicuna-7b, indicating the compressed virtual tokens can provide the original truncated information by recovering the cut-off parts after supervised fine-tuning.
Appendix B Results on BlueLM
We extra conduct experiments on BlueLM Team (2023) to verify the generality of UniICL. We demonstrate the result of understanding tasks in Table 5, of the generative tasks in Table 6.
Model #-shots CoLA-dev SST-2-dev IMDb Acc. BlueLM 0-shot 71.6 81.2 48.8 1-shot 69.6 82.6 64.8 2-shot 70.0 87.0 65.6 5-shot 70.5 88.6 68.7 UniICL 1-shot 69.6 81.2 65.4 2-shot 68.7 82.6 67.0 5-shot 71.7 87.0 70.4 UniICL 1-shot 69.8 80.0 62.0 2-shot 70.1 80.8 67.0 5-shot 71.8 85.6 69.6 8-shot 72.3 87.4 69.4 UniICL + $L_{ctr}$ 1-shot 70.1 80 69.6 2-shot 70.3 87.2 70.6 5-shot 71.1 89.2 71.0 8-shot 72.5 90.4 76.8
Table 5: The ICL results of understanding tasks with the backbone of BlueLM.
Method #-shots XSUM ARXIV R-1 R-2 R-L R-1 R-2 R-L BlueLM 0-shot 23.5 6.8 17.6 30.9 7.7 24.7 1-shot 19.1 4.8 12.1 23.0 3.6 19.0 UniICL 1-shot 24.0 6.9 18.0 31.4 7.7 25.2 2-shot 25.0 7.3 18.8 30.8 7.3 24.8 5-shot 25.0 7.3 18.8 30.8 7.3 24.8 UniICL 1-shot 25.2 7.4 18.9 31.6 7.9 25.4 2-shot 25.4 7.6 19.1 31.9 8.0 25.6 5-shot 26.5 7.9 20.3 32.1 8.0 25.5
Table 6: The ICL results of generative tasks with the backbone of BlueLM.
Appendix C Datasets & Metrics
C.1 Datasets
Dataset # words (96,512] (512,1024] (1024,1536] XSUM - 10,000 4,697 CICERO 10,000 - - XSUM (Ctr) 5,000
Table 7: The composition training set of UniICL. (m,n] represents the range of the number of words in each instance. XSUM (Ctr) is used for the second phase training in Equation 7.
Dataset Task In-Domain # Test # Demonstrations MS MARCO-dev Passage Ranking 6,980 - XSUM Text Summarization 1,500 204,045/20 ARXIV Text Summarization 1,500 203,037/20 CoLA-dev Lingustic Acceptability 1,041 67,349/20 SST2-dev Sentiment Classification 872 8,551/20 IMDb Sentiment Classification 1,500 25,000/20
Table 8: The details of involved evaluation datasets. -dev represents employing development set due to their test sets are inaccessible. # Demonstrations represent the number of demonstrations to be selected in high/low-resource ICL settings.
We mix up two public datasets for compression and selection augmentation training, described in Table 7. Additionally, UniICL achieves outstanding performance on out-domain evaluation, involving text summarization Narayan et al. (2018), passage ranking Nguyen et al. (2016), sentiment classification Maas et al. (2011); Socher et al. (2013), and linguistic acceptability Warstadt et al. (2018), more details referring to Table 8. UniICL selects demonstrations from its training set in high-resource ICL, and we fixed the number of candidate demonstrations to 20 for low-resource ICL evaluation.
C.2 Evaluation Metrics
ROUGE Lin (2004) is a widely adopted metric in many generative tasks that evaluate how similar the generated hypothesis is to the golden label. Therefore, ROUGE is used in our experiments to evaluate the quality responses generated conditioned on compressed virtual tokens, and we report the F-1 scores of ROUGE-1, ROUGE-2, and ROUGE-L (abbreviated R-1, R-2, R-L in the following), and we employed the files2rouge  [^10] https://github.com/pltrdy/files2rouge.
library in practice. Following the previous works, we report the accuracy of close-ended evaluation and MRR@10 for passage ranking.

### v2

@@
 Abstract
-In-context learning (ICL) facilitates large language models (LLMs) exhibiting spectacular emergent capabilities in various scenarios.
+In-context learning (ICL) enables LLMs to exhibit spectacular emergent capabilities in various scenarios.
 Unfortunately, introducing demonstrations easily makes the prompt length explode, bringing a significant burden to hardware.
@@
 In this paper, we propose an ICL framework UniICL, which Unifies demonstration selection and compression, and final response generation via a single frozen LLM.
-Specifically, UniICL first projects actual demonstrations and inference text inputs into short virtual tokens, respectively.
-Then, virtual tokens are applied to select suitable demonstrations by measuring semantic similarity within latent space among candidate demonstrations and inference input.
-Finally, inference text inputs together with selected virtual demonstrations are fed into the same frozen LLM for response generation.
-Notably, UniICL is a parameter-efficient framework that only contains 17M trainable parameters originating from the projection layer.
-We conduct experiments and analysis over in- and out-domain datasets of both generative and understanding tasks, encompassing ICL scenarios with plentiful and limited demonstration candidates.
-Results show that UniICL effectively unifies 12×12$\times$12 × compression, demonstration selection, and response generation, efficiently scaling up the baseline from 4-shot to 64-shot ICL in IMDb with 24 GB CUDA allocation [^1] The code and model will be released in the final version..
+UniICL leverages the understanding ability of well-trained LLMs to independently compress different demonstrations into compressed features, and then a learnable projection layer converts features to LLM-acceptable compressed virtual tokens.
+Apart from substituting original demonstrations to reduce input length, virtual tokens are again used to select potential demonstrations.
+Finally, current queries together with selected compressed virtual tokens are fed into the same frozen LLM for response generation.
+UniICL is a parameter-efficient framework that only contains 17M trainable parameters originating from the projection layer and a learnable embedding.
+We build UniICL upon two backbones and conduct experiments over in- and out-domain datasets of both generative and understanding tasks, encompassing ICL scenarios with plentiful and limited demonstration candidates.
+Results show that UniICL effectively unifies 12×12$\times$12 × compression, demonstration selection, and response generation, efficiently scaling up the baseline from 4-shot to 64-shot ICL with 24 GB CUDA allocation [^1] The code and model will be released in the final version..
 Unifying Demonstration Selection and Compression for In-Context Learning
-Jun Gao School of Computer Science and Technology Soochow University junegao1106@gmail.com
-Figure 1: The efficiency comparison between UniICL and other compression methods in IMDb with the number of shots increasing from 0 to 64.
-* and the break of line chart indicates occurring memory explodes.
-Figure 2: Compressing candidate demonstrations into spans of compressed virtual tokens and selecting one for the inference inputs.
-Compressed virtual tokens substitute the original demonstration together with the inference input, fed to the target LLM for prediction generation.
+Jun Gao1, Ziqiang Cao1∗, Wenjie Li2 School of Computer Science and Technology1 Department of Computer Science, Hong Kong Polytechnic University2
+Figure 1: UniICL compresses candidate demonstrations into spans of compressed virtual tokens and selects demonstrations similar to the query.
+Virtual tokens substitute the original demonstration together with the query, fed to the same LLM for prediction generation.
 1 Introduction
@@
 Despite its significant role in the era of LLMs, ICL also brings an enormous challenge to the input window.
-Specifically, inevitably introducing demonstrations directly causes length disaster Wang et al.
-(2024), which is more serious in long generative tasks such as question answering Ghosal et al.
-(2022) and text summarization Narayan et al.
+Specifically, inevitably introducing demonstrations directly causes length explosion Wang et al.
 (2024), bringing significant memory costs and decreasing inference throughput as described in Figure 8.
-Except for length disaster, Liu et al.
-(2021) points out that the quality of selected demonstrations significantly influences the ICL performance.
-The mentioned two issues derive two research lines focused on selecting salient demonstrations and utilizing demonstrations efficiently.
-Generally, researchers attempted to alleviate length explode Wingate et al.
+Except for length explosion, Liu et al.
+(2021) points out that the quality of the selected demonstrations significantly influences the ICL performance.
+Recent efforts in modifying model architecture significantly expand the input window Zheng et al.
+(2022); Wu et al.
+(2022); Ding et al.
+(2023); Bulatov et al.
+(2023).
+However, LLMs with million context windows still struggle to overcome performance degradation Liu et al.
+(2024).
+Researchers also attempted to alleviate length explosion via prompt pruning Jiang et al.
+(2023) or soft prompts Wingate et al.
 (2022); Mu et al.
 (2023); Ge et al.
-(2023) via soft prompts.
-The main idea was to train an independent compressor to compress the input into compact virtual tokens.
-Similarly, Liu et al.
+(2023).
+Their main idea is to train an independent compressor to compress the input into soft prompts.
+While in selecting powerful in-context demonstrations, Liu et al.
 (2021); Lu et al.
-(2021) scores each candidate demonstration for the given inference input via an extra ranker, and Ram et al.
+(2021) scores each candidate demonstration for the current queries via an extra ranker, and Ram et al.
 (2023); Wang et al.
 (2024) finds that a fine-tuned LLM is up to the task of ranking.
-However, the external compressor/ranker naturally leads to the Out-of-Distribution (OoD) problem between the compressor/ranker and the target LLM Jiang et al.
-(2023), requiring extra large-scale data to perform alignment.
-In addition, the standalone compressor or ranker incurs additional memory costs as it necessitates simultaneous loading alongside the target LLM.
-Motivated by the inherent semantic understanding ability of LLMs, we propose an ICL framework UniICL that unifies demonstration compression, demonstration selection, and response generation via the same frozen LLM.
-In UniICL, the same frozen LLM is applied thrice: it acts as the compressor to compress demonstrations into virtual tokens, as the selector to select suitable demonstrations based on the saliency measured by virtual tokens, and as the generator to produce final responses under the guidance of selected virtual demonstrations.
-Specifically, similar to the mask token in pretrained language understanding models Devlin et al.
-(2018); Liu et al.
-(2019), UniICL introduces the compression slot [M] and attaches $k\times[\text{M}]$ behind demonstrations to be condensed.
-The LLM forwards each processed demonstration and outputs corresponding $k$ hidden states on top of compression slots.
-Considering the latter compression slots naturally summarize the preceding demonstrations due to the transformer blocks, we train a simple projection layer to convert continuations to condensed virtual tokens.
-We briefly explain UniICL on the linguistic acceptability task concerning how it selects one demonstration among two compressed candidates and produces the prediction in Figure 2.
-On the one hand, virtual tokens effectively absorb semantic information from actual demonstrations, along with inference inputs, fed into the target LLM to perform generation as normal ICL.
-On the other hand, virtual tokens are again applied to measure saliency scores between demonstrations and given inference inputs within latent space, while demonstrations with higher scores are preferentially fed to the generator for response generation.
-Furthermore, UniICL compresses each demonstration independently of others, allowing caching virtual token sequences to configure a Demonstrations Bank (DB) for reusing, which will significantly bolster the inference throughput as demonstrated in Figure 1.
-Generally, since the compressor, selector, and generator are the same frozen LLM, UniICL merely equips 17M trainable parameters originating from a projection layer and one special embedding [M], and we collect 30k public data for training.
-We extensively evaluate UniICL on in- and out-domain datasets of both generative and understanding tasks, covering the mainstream high- and low-resource ICL [^2] High- and low-resource ICL are distinguished by the magnitude of candidates.
-The former usually selects demonstrations from plentiful candidates, and the latter merely selects from a handful of data..
-Results indicate that UniICL effectively substitutes 12$\times$ longer demonstrations with virtual tokens and scale up the inner LLM from 4-shot to 64-shot, significantly alleviating length disaster at an expense of 8% extra inference latency (without caching).
-Meanwhile, compared with the popular dense retriever, UniICL selects more suitable demonstrations in different tasks.
+However, the standalone compressor or ranker inevitably incurs additional memory costs as it necessitates simultaneous loading alongside the target LLM.
+To tackle these challenges, we propose a framework UniICL as illustrated in Figure 1, which leverages the semantic understanding ability of target LLMs developed during pre-training to compress and select demonstrations.
+UniICL keeps the target LLM frozen during training to avoid catastrophic forgetting and reduce training costs.
+Considering the frozen LLM within UniICL function as both compressor and generator, UniICL is training efficiently and does not need to load extra compression modules, significantly reducing memory costs for both training and inference.
+As the underlying LLMs in visual language models fail to understand visual features without an adapter, UniICL still requires converting compressed features into LLM-acceptable compressed virtual tokens.
+Apart from substituting lengthy demonstrations, virtual tokens are dense representations of original demonstrations, which can be naturally further applied to select potential demonstrations by measuring similarity.
+In this paper, the connector is a projection layer supervised-tuned under the language modeling objective of the generator during learning compression and jointly optimized by language modeling and contrastive objectives He et al.
+(2020) during learning selection.
+UniICL notices the fact that demonstrations in the ICL are independent of each other, which is ignored by previous compression studies Mu et al.
+(2023); Wingate et al.
+(2022); Ge et al.
+(2023).
+Therefore, UniICL independently compresses demonstrations and proposes to configure Demonstration Bank (DB) to cache compressed virtual tokens for further reusing without requiring repeated compression for the same demonstration.
+We evaluate UniICL which only contains 17M trainable parameters on a scope of benchmarks involving linguistic acceptability, semantic classification, text summarization, and passage reranking, and UniICL achieves outstanding performances.
+Then, the in-domain experiments show that UniICL built upon two backbones (Vicuna and BlueLM) effectively substitutes the 12$\times$ demonstrations with soft prompts, easily scaling up the inner LLM from 4-shot to 64-shot.
 Our main contributions are as follows:
-  • To our knowledge, we are the first to propose an ICL framework that unifies compression, selection, and generation via a single frozen LLM, inherently bypassing the OoD problem.
+  • To our knowledge, we are the first to propose an ICL framework that unifies compression, selection, and generation via a single frozen LLM.
   • UniICL is a memory-friend framework that enables LLMs to perform large-shot ICL on consuming GPUs.
-  • Simple application of virtual tokens outperforms the popular dense retriever in selecting suitable demonstrations after jointly optimized by compression and selection augmentation loss.
+  • UniICL proposes to configure Demonstration Bank to enhance ICL efficiency, avoiding repeated compression for the same demonstration.
 2 Related Work
-In this section, we simply rethink the methods focus on demonstration (prompt) compression and document retrieval.
 2.1 Soft Prompt Compression
@@
 To tackle the generalization problem, Mu et al.
-(2023) proposed to learn a Llama-7b to compress instruction to virtual tokens, but only compress instruction was not powerful enough since the demonstrations were much longer in practical.
+(2023) proposed to learn a Llama-7b to compress instruction to virtual tokens, but only compress instruction was not powerful enough since the demonstrations were much longer in practice.
 To compress the demonstrations, Chevalier et al.
-(2023) proposed AutoCompressor to generate compressed virtual tokens based on a fine-tuned LLM Zhang et al.
+(2023) proposed AutoCompressor to recurrently generate compressed virtual tokens based on a fine-tuned Llama Zhang et al.
 (2022).
-They first randomized segmented the texts with thousands of words into model-accepted range and then recursively generated soft prompts for each segment.
-Similarly, Ge et al.
+However, AutoCompressor broke the independence of demonstrations, and the recurrent compression increased inference latency.
+Ge et al.
 (2023) proposed ICAE that employed a LoRA-adopted Llama-7b Touvron et al.
-(2023) to compress the processed demonstrations to compact virtual tokens, and then fed to a frozen LLM to perform generation.
-Unlike previous works, the underlying LLM in UniICL was kept frozen during training and was trained on only 30k data.
+(2023) to compress the processed demonstrations to compact virtual tokens, while ICAE still struggled to overcome quite long inputs.
 2.2 Extractive Compression
@@
 However, UniICL naturally bypassed the distribution-aligned module, since the compressor and the target LLM were the same.
+Methods Additional Compressor Compression Tool # Trainable Parameters Train Size LLMLingua YES Pruning 7B 57k AutoCompressor NO Soft Prompt 7B UNKNOWN ICAE YES Soft Prompt 70M 240k UniICL NO Soft Prompt 17M 30k
+Table 1: Comparsion among recent compression methods and UniICL. Compression Tool represents the involved compression technique of different methods.
 3 Methodology
 We propose UniICL, a parameter-efficient ICL framework that unifies demonstration compression, demonstration selection, and response generation via a single LLM.
-Specifically, UniICL applies a general frozen LLM thrice, as a compressor to compress demonstrations into short virtual tokens, as a selector to choose suitable demonstrations for ICL, and as a generator to produce responses.
-As for the selection of the underlying LLM, previous work has proposed that the Decoder-only model performs better than the Encoder-Decoder model in prompt compression Mu et al.
+As for the selection of the underlying LLM, previous work has proved that the Decoder-only model performs better than the Encoder-Decoder model in prompt compression Mu et al.
 (2023).
 We follow this conclusion and employ popular Vicuna-7b Zheng et al.
-(2023) and BlueLM-7B [^3] We mainly utilize Vicuna-7B to verify the effectiveness, the experiments of BlueLM will be exhibited in Appendiex B for concise.
-Team (2023) as the de facto backbone in UniICL.
-We present a comparison between UniICL and other recent compression methods in Table 1.
-Without regard to the Lora-adopted compressor, ICAE Ge et al.
-(2023) is most similar to UniICL in architecture and compression mode that parameter-efficiently compresses actual demonstrations into dense soft prompts.
+(2023) and BlueLM-7B [^2] We mainly discuss UniICL built on Vicuna-7B, the experiments of BlueLM will be exhibited in Appendiex B.
+Team (2023) as the underlying backbone in UniICL.
+Figure 2: Differences of compression methods in formulating compressed virtual tokens in ICL.
+We present a comparison of training costs between UniICL and other recent compression methods in Table 1.
+Additionally, we illustrate differences in formulating virtual tokens for compression methods based on the soft prompt in Figure 2.
+To explain plainly, we ideally assume the compressor within three compression methods based on soft prompts has the window limitation of $L$, and has the same compression ratio, ignoring the length of soft prompts.
+In the 2-shot scenario, demonstrations $D_{1}$ and $D_{2}$ have a length of $L$.
+Considering the AutoCompressor, the concatenated demonstrations will be divided back into two segments, and the AutoCompressor compresses each segment step-by-step, bringing two times non-parallel compression.
+When it comes to ICAE, merely $D_{1}$ is accessible for the compressor and others will be read by no means.
+AutoCompressor shows advantages in the readable prompt length, but is short in efficiency, while ICAE has a constant compression complexity but struggles to approach relatively long inputs.
+Additionally, AutoCompressor makes the compression of later demonstrations be conditioned on the previous ones, which breaks demonstration independence in ICL.
+Combining the advantages of AutoCompressor and ICAE, UniICL compresses demonstrations independently and introduces virtual tokens concatenation to overcome long prompt challenges.
+In the $N$-shot settings, the number of practical compression steps can be calculated as $\lceil\frac{N}{k}\rceil$, where $k$ indicates that a single GPU is capable of compressing $k$ demonstrations in a batch.
+When the GPU capacity is sufficient, k equals $N$, which is the scenario of ICAE that compresses all segments in a time but UniICL drops nothing, while it degenerates to the AutoCompressor scenario that compresses segments step-by-step, when the GPU capacity is only sufficient to set $k=1$.
 Figure 3: Demonstration compression.
-Methods Additional Compressor Compression Tool # Trainable Parameters Train Size LLMLingua YES Pruning 7B 57k AutoCompressor NO Soft Prompt 7B UNKNOWN ICAE YES Soft Prompt 70M 240k UniICL NO Soft Prompt 17M 30k
-Table 1: Comparsion among recent compression methods and UniICL. Compression Tool represents the involved compression technique of different methods.
-3.1 Preliminary
-Empirically, a prompt may encompass in-context demonstrations and an inference input in the ICL scenario.
-To avoid misleading and follow the existing definitions, we assume a prompt $P=(x^{D}_{1},x^{D}_{2},...,x^{D}_{n},x^{I})$, where $x^{D}_{i}$ is the $i$th demonstrations consist of an input-output pair (e.g.
-[DOC] -> [Sum]), and $x^{I}$ is the inference input (e.g.
-documents in the summarization task).
-We first initialize the compression slots [M] $\in\text{R}^{d}$ from the rarely used embedding of corresponding LLM, with gradient updating during training, $d$ is the dimension of the input embeddings.
-The compression ratio $r$ is a hyper-parameter, which is dynamically sampled from 2 to 16 during training, to calculate the specific number of compressed slots by $k=|x^{D}|/{r}$.
 3.1 Demonstration Compression
-For each demonstration in $P$, UniICL activates the compression slots to aggregate information from demonstrations in the forward propagation of frozen Vicuna respectively, as illustrated in Figure 3.
-Practically, we first attach $k$ compression slots $=k\times\text{[M]}$M] to each demonstration $x^{D}_{i}$, formatting modified prompt fed to the Vicuna.
-Then, frozen Vicuna forwards the modified prompts and outputs the last hidden states $H_{i}=(h_{1},h_{2},...,h_{k})$ corresponding to $k$ compression slot, dropping the others444$\oplus$ means token-level concatenation.:
-_$\_,H_{i}=\text{Encode}(x^{D}_{i}\oplus M)$ (1)
-Due to the transformer block in Vicuna, $H_{i}$ is compelled to attend to the preceding actual tokens and UniICL inserts a linear layer to project $H_{i}$ into compressed virtual tokens $C^{i}=(c_{1},c_{2},...,c_{k})$:
-$c_{j}=W_{p}\cdot h_{j},$ (2)
+UniICL introduces memory slots [M], a learnable embedding initialized from a rarely used embedding of the target LLM.
+UniICL activates the compression slots to absorb information from demonstrations in the forward propagation of frozen Vicuna, as illustrated in Figure 3.
+We first attach $k$ compression slots $=k\times\text{[M]}$M] to each demonstration $D_{i}$, formatting modified prompt fed to the Vicuna.
+Then, frozen Vicuna forwards the modified prompts and outputs the last hidden states $H_{i}=(h_{1},h_{2},...,h_{k})$ on top of the $k$ compression slots, dropping the others333$\oplus$ means token-level concatenation.:
+_$\_,H^{i}=\mathrm{forward}(D_{i}\oplus M)$ (1)
+Due to the attention mechanism, $H^{i}$ is compelled to attend to the preceding actual tokens.
+Then, UniICL inserts a projection layer to convert $H^{i}$ into LLM-acceptable compressed virtual tokens $C^{i}=(c^{i}_{1},c^{i}_{2},...,c^{i}_{k})$:
+$c^{i}_{j}=W_{p}\cdot h^{i}_{j},$ (2)
 where $W_{p}$ is the parameters of the projection layer.
@@
 3.2 Demonstration Selection
-Except for substituting the origin demonstrations for generation, activated virtual tokens $C^{i}$ are again applied for demonstration selection among candidates, as illustrated in Figure 4.
-Specifically, given an inference input $x^{I}$ and its candidate demonstrations $(x^{D}_{1},x^{D}_{2},...,x^{D}_{n})$, UniICL obtains their latent representation after average pooling:
-$\bar{C^{i}}=\frac{1}{k}\sum_{j=1}^{k}c_{j},$ (3)
-denote as $q$ for the input and $(k_{1},k_{2},...,k_{n})$ for candidate demonstrations for simlifying.
-We define the cosine similarity between $q$ and $k_{i}$ is the $i$th demonstration saliency score $S_{i}$:
-$S_{i}=cos(q,k_{i}).$ (4)
+Except for substituting the origin demonstrations for generation, compressed virtual tokens $C^{i}$ are representative of demonstrations that can be again applied for demonstration selection, as illustrated in Figure 4.
+Specifically, given a query $Q$ and its candidate demonstrations $(D_{1},D_{2},...,D_{n})$, UniICL obtains their latent representation after average pooling:
+$\bar{C^{i}}_{Q/D}=\frac{1}{k}\sum_{j=1}^{k}c_{j}.$ (3)
+We define the $i$-th demonstration saliency score $S_{i}$ via the cosine similarity between Query and demonstrations:
+$S_{i}=\mathrm{cos}(\bar{C_{Q}},\bar{C}_{D}^{i}).$ (4)
+Demonstrations are then reranked according to their saliency scores.
 3.3 In-context Generation
-We employ the frozen Vicuna again to generate responses with the guiding of concatenated virtual demonstrations, as illustrated in Figure 5.
-For $m$-shot in-context learning, we obtain $m$ virtual demonstrations after previously mentioned slot activating and demonstration selecting, denoted as $C^{1}$ to $C^{m}$.
+We employ the frozen Vicuna again to generate responses with the guiding of concatenated virtual demonstrations and queries, as illustrated in Figure 5.
+For $m$-shot in-context learning, we obtain $m$ virtual tokens after demonstration compression and selection, denoted as $C^{1}$ to $C^{m}$.
 Then, we horizontally concatenate them, keeping their relative position unmodified.
-Finally, the concatenated virtual demonstrations together with actual inference inputs are fed into Vicuna, performing auto-regressive generation as normal:
-$y_{i}=\text{Decode}(C^{1},...,C^{m};x^{I};y_{<i})$ (5)
-Except for the generative manner, virtual tokens are conveniently transferred to close-ended evaluation without modifications for understanding tasks through providing the candidate answers and measuring their perplexity (PPL)  [^5] https://huggingface.co/docs/transformers/perplexity respectively, e.g.
+Finally, the concatenated virtual tokens together with actual inference inputs are fed into Vicuna, performing auto-regressive generation as normal:
+$y_{i}=\mathrm{generate}(C^{1},...,C^{m};Q;y_{<i})$ (5)
+Except for the generative manner, virtual tokens are conveniently transferred to close-ended evaluation for understanding tasks through providing the candidate answers and measuring the label space PPL  [^4] https://huggingface.co/docs/transformers/perplexity, e.g.
 $(ppl^{+},ppl^{-}$) for sentiment classification:
-$\text{Prediction}=argmin(ppl^{+},ppl^{-}),$ (6)
+$y=\mathrm{argmin}(ppl^{+},ppl^{-}),$ (6)
 where answers with PPL closest to 1 are judged to be the current prediction.
@@
 The trainable parameters in UniICL are merely 17M originating from the projection layer $W_{p}$ and the introduced compression slot [M].
-We first use the common language modeling loss to optimize UniICL and get a base compression model.
-Then we add InfoNCE He et al.
-(2020) to augment the demonstration selection ability:
+The projection layer is optimized with the language modeling objective of Vicuna to learn a base compression model.
+Then InfoNCE He et al.
+(2020) joint with language modeling objective are used to augment the demonstration selection ability of the base compression model:
 $\mathcal{L}=\mathcal{L}_{lm}+\mathcal{L}_{ctr}.$ (7)
-Specifically, during the first training phase, we slice the source input of each training instance into two parts and randomly compress one, denoted the compression part as $x_{c}$ and the unmodified as $x_{u}$, respectively.
-Afterward, we attach $M$ to $x_{c}$ and get virtual tokens $C$ on top of the compression slot sequence, as described in Equ.
+Specifically, we slice the source input of each training instance into two parts and randomly compress one, denoted the compression part as $x_{c}$ and the unmodified as $x_{u}$.
+Afterward, we attach $M$ to $x_{c}$ and get virtual tokens $C$ on top of the compression slots, as described in Equ.
 1 and Equ.
 2.
-Therefore, the language modeling loss $\mathcal{L}_{lm}$ if obtained as:
-$\mathcal{L}_{lm}=-\frac{1}{|y|}\sum_{t=0}logP(y_{t}|x_{2};C;y_{<t}),$ (8)
-where $y=(y_{1},y_{2},...)$ is the reference label of the current training instasnce.
-Notably, distinguished from ICAE which always puts soft prompts in the front, UniICL keeps the relative position of $C$ and $x_{u}$ unchanged to make compressed tokens insensitive to their position in prompts.
-Namely, compressed tokens will be placed in the former (latter) if $x_{c}$ is in front (behind) of $x_{u}$ in the original sequence.
-Additionally, aiming to concatenate different virtual token sequences from different demonstrations for practical employment of ICL, we introduce concatenation compression that evenly slices the extra-long compression parts into two sub-segments during training, compresses them separately, and concatenates their virtual tokens horizontally.
-Consequently, we utilize contrastive learning for selection augmentation and mine positives $x^{D}_{+}$ and negatives $x^{D}_{-}$ as illustrated in Figure 6.
-Specifically, given each training instance $x^{I}$ and n candidate demonstrations $(x^{D}_{1},x^{D}_{2},...,x^{D}_{n})$ from two non-crossing training subsets, we employ Vicuna to calculate the PPL concerning the golden label of $x^{I}$, denoted as $ppl^{I}$.
-Then, we provide the $i$th demonstration and calculate its PPL, denoted as $(ppl^{D}_{i},i\in[1,n])$.
-We count $ppl^{I}$ as the baseline and calculate candidate relative PPL gains:
-$\widetilde{ppl^{D}_{i}}=ppl^{I}-ppl^{D}_{i},i\in[1,n].$ (9)
-Hence, $k^{+}$ $(k^{-})$ is the latent representation of demonstrations that furthest reduces (increases) $ppl^{I}$ as processed in Equation 3, and the contrastive loss $\mathcal{L}_{ctr}$ can be formulated as:
-$\mathcal{L}_{ctr}=\frac{exp(cos(q,k^{+}))}{exp(cos(q,k^{+}))+exp(cos(q,k^{-}))}.$ (10)
+Therefore, the language modeling loss $\mathcal{L}_{lm}$ is obtained as:
+$\mathcal{L}_{lm}=-\frac{1}{|y|}\sum_{t=0}logP(y_{t}|Q;C;y_{<t}),$ (8)
+where $y$ is the reference label of the current training instance.
+Additionally, to approach the large-shot settings without significant truncation, we introduce concatenation compression.
+When $x_{c}$ exceeds the window limitation for compression, UniICL further divides $x_{c}$ into acceptable ranges and compresses them independently to get local virtual tokens.
+Then, these virtual tokens of several segments will be concatenated to formulate global virtual tokens to replace $x_{c}$.
+Consequently, we utilize contrastive learning for selection augmentation and mine positives and negatives as illustrated in Figure 6.
+Specifically, given each training instance $Q$ and $n$ candidate demonstrations $(D_{1},D_{2},...,D_{n})$ from two non-crossing training subsets, we employ Vicuna to calculate the PPL concerning the golden label of $Q$, denoted as $ppl^{Q}$.
+Then, we provide the $i$-th demonstration and calculate PPL concerning the golden label of $Q$, denoted as $(ppl^{D}_{i},i\in[1,n])$.
+We count $ppl^{Q}$ as the baseline and calculate candidate relative PPL gains:
+$\widetilde{ppl^{D}_{i}}=ppl^{Q}-ppl^{D}_{i},i\in[1,n].$ (9)
+Hence, $C_{D}^{+}$ $(C_{D}^{-})$ is the representation of demonstrations $D^{+}$ ($D^{-}$) that furthest reduces (increases) $ppl^{Q}$ as processed in Equ.
+3, and the contrastive loss $\mathcal{L}_{ctr}$ can be formulated as:
+$\mathcal{L}_{ctr}=\frac{exp(cos(C_{Q},C_{D}^{+}))}{exp(cos(C_{Q},C_{D}^{+}))+exp(cos(C_{Q},C_{D}^{-}))}.$ (10)
 In particular, if all relative PPL gains are less than 0, namely none of the candidate demonstrations help guide Vicuna to generate the golden label, we will apply the other set of candidates.
-Notably, the inner compression ratio $r$ for each training step is dynamic and randomly sampled from [2, 16] during training.
-After the two training phases, we blend 2,000 instances from the in-domain validation sets to evaluate the quality of virtual demonstrations provided to LLMs under various ratios.
-Ultimately, the preferred scheme involves compressing the demonstrations with $r=12$, more details refer to Figure 7.
 Figure 6: Negatives mining pipeline.
 4 Experiment
-In this section, we comprehensively evaluate UniICL over out-domain datasets to verify its effectiveness.
-Additionally, we detailedly introduce of involved datasets and evaluation metrics in Appendiex C.
 4.1 Baselines
@@
 ICAE [^7] https://github.com/getao/icae.
-compresses demonstrations into soft prompts via a LoRA-adapted Llama2-7b, which is most similar to UniICL in model architecture and compression mode.
+compresses demonstrations into soft prompts via a LoRA-adapted Llama2-7b.
 Additionally, since selection augmentation is involved in the training of UniICL, we utilize the popular Sentence-BERT (S-BERT) Reimers and Gurevych (2019) as the dense retriever to construct an ICL pipeline for the above methods, serving as simple but effective selection-based baselines.
 4.2 Settings
-Considering the involved datasets and computation efficiency, we set the max allowed input length limit to 1,024 for both compression and generation.
-For a fair comparison, we set the allowed window of baselines to 1,024 and their compression ratio to 12.
+Considering the involved datasets and computation efficiency, we set the max allowed input length limit to 512 for both compression and generation.
+For a fair comparison, we set the allowed window of baselines to 512 and their compression ratio to the same as UniICL.
 We fix the learning rate to 8e-5 and use Adam as the optimizer, and the effective batch size is 32 (8 GPUs data parallelism and 4 steps gradient accumulation).
-Additionally, we conducted all experiments on 8*NVIDIA A5000 24G GPUs based on BFloat 16 data type, and we set the evaluated shot to 8 for understanding tasks and 5 for generative tasks for illustration.
-Except for the MS MARCO preprocessed by Liang Wang et al.
-(2022)  [^9] https://github.com/microsoft/unilm/tree/master/simlm., the other evaluation datasets are publicly achieved at hugging face.
-We apply S-BERT to pre-rank and output the top 10 similar candidates from training sets according to each inference input and employ UniICL to perform selection among them in practice due to computation efficiency for high-resource ICL.
-On the contrary, the low-resource ICL setting fixes the candidate demonstrations to 20 for all inference inputs, performing pre-ranking and selecting as well.
+Additionally, we conducted all experiments on 8*NVIDIA A5000 24G GPUs based on BFloat 16 data type, and we set the evaluated shot to 8 for understanding tasks and 5 for generative tasks for illustration because of marginal ICL gains and memory costs.
+We apply S-BERT to pre-rank and output the top 10 similar candidates from training sets according to each inference input for all baselines.
+UniICL is employed to perform selection among them in practice due to computation efficiency for high-resource ICL.
+On the contrary, the low-resource ICL setting fixes the random candidate demonstrations to 20 for all inference inputs, performing pre-ranking and selecting as well.
 4.3 Results
@@
 AutoCompressor 29.3
-LLMLingua 27.8
 ICAE 30.2
@@
 Notably, we don’t compare UniICL with the popular retrieval models Reimers and Gurevych (2019); Wang et al.
-(2024) since most of them are fine-tuned on this dataset, unfair for comparison.
+(2024) since most of them are fine-tuned on this dataset, which is unfair for comparison.
 Figure 7: The overall sensitivity analysis of compression ratio.
@@
 Therefore, we set the compression ratio to 12 by default and apply this ratio to all experiments.
-The 1024×1024$\times$1024 × compression ratio is equal to compressing anything to a single virtual token, due to the maximum allowed input length for compression being 1,024.
+The 512×512$\times$512 × compression ratio is equal to compressing anything to a single virtual token, due to the maximum allowed input length for compression being 512.
+Figure 8: The efficiency comparison between UniICL and other compression methods in CoLA with the number of shots increasing from 0 to 64.
+Memory explodes are represented as *, corresponding to the break of the line chart.
+Method GPUHours TFLOPs TMACs Vicuna 1.5 86,20 4,309 Vicuna-1k 1.9 31,664 15,832 UniICL 1.6 22,437 11,218
+Table 4: The computation efficiency of UniICL.
 5.2 Efficiency Analysis
-In UniICL, we incorporate an additional 17M trainable parameters into the 7b backbone, accounting for an approximate increase of 2.4%.
+In UniICL, we incorporate an additional 17M trainable parameters into the 7b backbone, accounting for an approximate increase of 0.24%.
 We additionally evaluate the memory costs inference latency of UniICL and other compression methods in Figure 8.
-Notably, compressing each demonstration independently and allowing horizon concatenation among different virtual demonstrations convert the extra compression latency be a one-shot deal that the virtual tokens can be stored, formulating a Demonstration Bank (DB) for further reusing.
-In this case, UniICL will eliminate the extra latency if the selected demonstrations have been compressed and cached (UniICL+Caching).
+With the help of the Demonstration Bank (DB), UniICL will eliminate the extra latency if the selected demonstrations have been compressed and cached (UniICL+Caching).
 Despite this, parallel computation facilitates the compressing process, resulting in minimal throughput degradation (UniICL and Baseline).
-What’s more, the naive 7B LLM occurs memory explosion for 8-shot settings, while UniICL successfully scales up to 64-shots within 24GB CUDA allocation.
-5.3 Necessarity of UniICL
-Intuitively, UniICL applies a “heavy” LLM as the compressor, but the generator and the compressor share all parameters without loading double parameters as LLMLingua, bypassing the OoD problem between the compressor and generator.
-Additionally, ICAE switches the LoRA parameters for compression and generations respectively, which appear to make the compressor and the generator share parameters.
-However, the compressor and the generator in ICAE are still two different LLMs inherently after assembling LoRA weight.
-Furthermore, the continual switch also breaks the latency-free merit of LoRA, resulting in more inference latency than the projection layer in UniICL, as depicted in Figure 1.
+What’s more, the naive 7B LLM occurs memory explosion for 8-shot settings and other compression methods perform up to 32-shot, while UniICL successfully scales up to 64-shots within 24GB CUDA allocation.
+Additionally, We demonstrate the inference computation and GPU hours in Table 4, by using 1,024 random legal tokens as inputs and forcing models to generate 128 tokens.
+Notably, UniICL (without DB) compresses the former half, and the latter half is fed into the generator directly, while Vicuna and Vicuna-1k are distinguished in window limitations.
+Results indicate that minimal GPU hours increased due to the parallel computation of forward, although the extra compression of UniICL surges the computation.
+Additionally, Vicuna with a 1k window limitation surges both GPU hours and TFLOPs because long input brings significant computation and latency in generation.
+5.3 Training Deviation
+To quantify the performance gains brought by the learnable projection layer.
+We tune Vicuna and BlueLM with comparable parameters (17M) with LoRA in Table 5, setting the rank to 32.
+UniICL still outperforms LoRA-adapted LLMs with a 512 window limitation, indicating that the truncation indeed brings performance degradation.
 6 Conclusion
@@
 Appendix A In-Domain Evaluation
-Method XSUM CICERO R-1 R-2 R-L R-1 R-2 R-L Vicuna 19.9 5.0 13.5 17.3 3.3 14.3    +UniICL 27.2 9.2 22.5 24.6 5.1 21.7
-Table 4: The in-domain results of XSUM and CICERO. Vicuna generates responses conditioned on the truncated inference input and UniICL first compresses the over-length input into virtual tokens.
-We conduct the zero-shot in-domain generation evaluation on the entire test set of XSUM and CICERO in Table 4, by compressing the latter half to virtual tokens and keeping the former unmodified.
-UniICL significantly outperforms the baseline Vicuna-7b, indicating the compressed virtual tokens can provide the original truncated information by recovering the cut-off parts after supervised fine-tuning.
+Backbone Method XSUM CICERO R-1 R-2 R-L R-1 R-2 R-L Vicuna-7b Vicuna 19.9 5.0 13.5 17.3 3.3 14.3    +LoRA 25.4 7.5 17.3 28.1 10.5 25.6 Vicuna-1k 27.3 8.7 19.7 30.5 11.3 27.4    +LoRA 31.2 11.0 23.1 34.1 13.5 30.2 UniICL 30.0 10.2 22.3 32.6 12.2 28.8 BlueLM-7b BlueLM 15.0 3.6 10.4 17.6 3.1 15.0    +LoRA 23.1 7.6 17.4 21.9 7.8 19.8 BlueLM-1k 28.1 9.9 22.8 25.1 9.2 23.1    +LoRA 30.8 10.5 24.6 31.2 10.8 27.4 UniICL 30.4 10.2 23.7 29.2 10.0 26.6
+Table 5: The in-domain results and ablation studies on XSUM and CICERO. 1k represents the extending 1k window limitation, while others have a limitation of 512.
+We conduct the zero-shot in-domain generation evaluation on the entire test set of XSUM and CICERO in Table 5 by compressing the latter half to virtual tokens and keeping the former unmodified.
+UniICL significantly outperforms the baselines, indicating the compressed virtual tokens can provide the original truncated information by recovering the cut-off parts after supervised fine-tuning.
+Although extending the window to 1k, Vicuna and BlueLM still underperform UniICL, indicating that compressed virtual tokens filter noise information to some extent.
 Appendix B Results on BlueLM
@@
 Table 6: The ICL results of understanding tasks with the backbone of BlueLM.
-Method #-shots XSUM ARXIV R-1 R-2 R-L R-1 R-2 R-L BlueLM 0-shot 23.5 6.8 17.6 30.9 7.7 24.7 1-shot 19.1 4.8 12.1 23.0 3.6 19.0 UniICL 1-shot 24.0 6.9 18.0 31.4 7.7 25.2 2-shot 25.0 7.3 18.8 30.8 7.3 24.8 5-shot 25.0 7.3 18.8 30.8 7.3 24.8 UniICL 1-shot 25.2 7.4 18.9 31.6 7.9 25.4 2-shot 25.4 7.6 19.1 31.9 8.0 25.6 5-shot 26.5 7.9 20.3 32.1 8.0 25.5
+Method #-shots XSUM ARXIV R-1 R-2 R-L R-1 R-2 R-L BlueLM 0-shot 15.0 3.6 10.4 30.9 7.7 24.7 1-shot 19.1 4.8 12.1 23.0 3.6 19.0 UniICL 1-shot 24.0 6.9 18.0 31.4 7.7 25.2 2-shot 25.0 7.3 18.8 30.8 7.3 24.8 5-shot 25.0 7.3 18.8 31.9 7.8 26.0 UniICL 1-shot 25.2 7.4 18.9 31.6 7.9 25.4 2-shot 25.4 7.6 19.1 31.9 8.0 25.6 5-shot 26.5 7.9 20.3 32.1 8.0 25.5
 Table 7: The ICL results of generative tasks with the backbone of BlueLM.
@@
 C.1 Datasets
-Dataset # words (96,512] (512,1024] (1024,1536] XSUM - 10,000 4,697 CICERO 10,000 - - XSUM (Ctr) 5,000
+Dataset # words (96,512] (512,1024] (1024,1536] XSUM - 10,000 4,697 CICERO 10,000 - - SUPER-NI - 10,000 7,000 XSUM (Ctr) 5,000
 Table 8: The composition training set of UniICL. (m,n] represents the range of the number of words in each instance. XSUM (Ctr) is used for the second phase training in Equation 7.

### v3

@@
-Unifying Demonstration Selection and Compression for In-Context Learning
+UniICL: An Efficient Unified Framework Unifying Compression, Selection, and Generation
 Abstract
-In-context learning (ICL) enables LLMs to exhibit spectacular emergent capabilities in various scenarios.
-Unfortunately, introducing demonstrations easily makes the prompt length explode, bringing a significant burden to hardware.
-In addition, random demonstrations usually achieve limited improvements in ICL, necessitating demonstration selection among accessible candidates.
-Previous studies introduce extra modules to perform demonstration compression or selection independently.
-In this paper, we propose an ICL framework UniICL, which Unifies demonstration selection and compression, and final response generation via a single frozen LLM.
-UniICL leverages the understanding ability of well-trained LLMs to independently compress different demonstrations into compressed features, and then a learnable projection layer converts features to LLM-acceptable compressed virtual tokens.
-Apart from substituting original demonstrations to reduce input length, virtual tokens are again used to select potential demonstrations.
-Finally, current queries together with selected compressed virtual tokens are fed into the same frozen LLM for response generation.
-UniICL is a parameter-efficient framework that only contains 17M trainable parameters originating from the projection layer and a learnable embedding.
-We build UniICL upon two backbones and conduct experiments over in- and out-domain datasets of both generative and understanding tasks, encompassing ICL scenarios with plentiful and limited demonstration candidates.
-Results show that UniICL effectively unifies 12×12$\times$12 × compression, demonstration selection, and response generation, efficiently scaling up the baseline from 4-shot to 64-shot ICL with 24 GB CUDA allocation [^1] The code and model will be released in the final version..
-Unifying Demonstration Selection and Compression for In-Context Learning
-Jun Gao1, Ziqiang Cao1∗, Wenjie Li2 School of Computer Science and Technology1 Department of Computer Science, Hong Kong Polytechnic University2
-Figure 1: UniICL compresses candidate demonstrations into spans of compressed virtual tokens and selects demonstrations similar to the query.
-Virtual tokens substitute the original demonstration together with the query, fed to the same LLM for prediction generation.
+In-context learning (ICL) enhances the reasoning abilities of Large Language Models (LLMs) by prepending a few demonstrations.
+It motivates researchers to introduce more examples to provide additional contextual information for the generation.
+However, existing methods show a significant limitation due to the problem of excessive growth in context length, which causes a large hardware burden.
+In addition, shallow-relevant examples selected by off-the-shelf tools hinder LLMs from capturing useful contextual information for generation.
+In this paper, we propose UniICL, a novel Unified ICL framework that unifies demonstration compression, demonstration selection, and final response generation.
+Furthermore, to boost inference efficiency, we design a tailored compression strategy that allows UniICL to cache compression results into Demonstration Bank (DB), which avoids repeated compression of the same demonstration.
+Extensive out-of-domain evaluations prove the advantages of UniICL in both effectiveness and efficiency.
+UniICL: An Efficient Unified Framework Unifying Compression, Selection, and Generation
+[…]
+Figure 1: (a) Prompt compression methods that indiscriminately compress both demonstrations and queries.(b) Retrieval-based demonstration selection methods select lexical demonstrations.
+(c) UniICL discriminately compresses demonstrations and performs selection upon the compression results.
 1 Introduction
@@
 (2021); Wang et al.
-(2023b) exhibits powerful performance in practical applications with the emergence of scaled-up Transformer-based Vaswani et al.
-(2017) Large Language Models (LLMs) Wang et al.
-(2023c); Yang et al.
+(2023b) to enhance the reasoning ability of Large Language Models (LLMs) with a few demonstrations prepended Wang et al.
+(2023d); Yang et al.
 (2023); Wei et al.
@@
 (2022).
-In ICL, the provided demonstrations activate the pre-training knowledge of LLMs to allow them to perform well on various downstream tasks such as text summarization Wang et al.
-(2023c); Yang et al.
-(2023) and text classification Min et al.
-(2022) without gradient updating of billion parameters.
-Despite its significant role in the era of LLMs, ICL also brings an enormous challenge to the input window.
-Specifically, inevitably introducing demonstrations directly causes length explosion Wang et al.
-(2024), bringing significant memory costs and decreasing inference throughput as described in Figure 8.
-Except for length explosion, Liu et al.
-(2021) points out that the quality of the selected demonstrations significantly influences the ICL performance.
-Recent efforts in modifying model architecture significantly expand the input window Zheng et al.
+Inspired by its outstanding performance, researchers explored applying ICL on many tasks such as text summarization Wang et al.
+(2023d); Yang et al.
+(2023); Gao et al.
+(2024a), sentiment classification, and linguistic acceptability Min et al.
+(2022); Wang et al.
+(2019).
+However, two challenges hinder the impact of ICL currently: (1) concatenated demonstrations directly surge the input length, causing a large hardware burden; (2) the prepended demonstrations are randomly sampled or selected via off-the-shelf tools which tend to provide shallow relevant demonstrations, hindering LLMs from capturing useful contextual information for generation.
+Existing work tackles the two challenges separately.
+To alleviate input length surge, on the one hand, many efforts are made in modifying model architecture to accommodate longer contexts Zheng et al.
 (2022); Wu et al.
@@
 (2023).
-However, LLMs with million context windows still struggle to overcome performance degradation Liu et al.
+These methods usually require training models from scratch, and models with a million context windows still struggle to overcome performance degradation Liu et al.
 (2024).
-Researchers also attempted to alleviate length explosion via prompt pruning Jiang et al.
-(2023) or soft prompts Wingate et al.
+On the other hand, recent studies attempt to shorten inputs through prompt compression Wingate et al.
 (2022); Mu et al.
+(2023); Jiang et al.
 (2023); Ge et al.
+(2023); Gao et al.
+(2024b).
+However, these compression methods are not applicable to ICL because they indiscriminately compress both demonstrations and queries into virtual tokens.
+For instance, as illustrated in Fig.
+1(a), the task entails justifying whether the query is grammatically acceptable.
+The latter generator makes responses only according to virtual tokens generated by the compressor, resulting in a wrong answer [^1] I hope to would study in Facnce (France).
+More importantly, current compression methods are costly to train Wingate et al.
+(2022); Mu et al.
+(2023); Jiang et al.
+(2023), and compressors are either limited to compressing within the original model’s allowed input length Mu et al.
+(2023); Jiang et al.
+(2023); Ge et al.
+(2023) or bringing significant inference latency Wingate et al.
 (2022).
-Their main idea is to train an independent compressor to compress the input into soft prompts.
-While in selecting powerful in-context demonstrations, Liu et al.
-(2021); Lu et al.
-(2021) scores each candidate demonstration for the current queries via an extra ranker, and Ram et al.
+Retrieval-based In-context Example Selection (RICES) methods Alayrac et al.
+(2022) integrate an off-the-shelf pre-training model to select demonstrations similar to the queries at a shallow level.
+These demonstrations usually contain redundant information and bring minimal benefits for the final generation Liu et al.
+(2021); Ram et al.
 (2023); Wang et al.
-(2024) finds that a fine-tuned LLM is up to the task of ranking.
-However, the standalone compressor or ranker inevitably incurs additional memory costs as it necessitates simultaneous loading alongside the target LLM.
-To tackle these challenges, we propose a framework UniICL as illustrated in Figure 1, which leverages the semantic understanding ability of target LLMs developed during pre-training to compress and select demonstrations.
-UniICL keeps the target LLM frozen during training to avoid catastrophic forgetting and reduce training costs.
-Considering the frozen LLM within UniICL function as both compressor and generator, UniICL is training efficiently and does not need to load extra compression modules, significantly reducing memory costs for both training and inference.
-As the underlying LLMs in visual language models fail to understand visual features without an adapter, UniICL still requires converting compressed features into LLM-acceptable compressed virtual tokens.
-Apart from substituting lengthy demonstrations, virtual tokens are dense representations of original demonstrations, which can be naturally further applied to select potential demonstrations by measuring similarity.
-In this paper, the connector is a projection layer supervised-tuned under the language modeling objective of the generator during learning compression and jointly optimized by language modeling and contrastive objectives He et al.
-(2020) during learning selection.
-UniICL notices the fact that demonstrations in the ICL are independent of each other, which is ignored by previous compression studies Mu et al.
-(2023); Wingate et al.
-(2022); Ge et al.
 (2024).
-Therefore, UniICL independently compresses demonstrations and proposes to configure Demonstration Bank (DB) to cache compressed virtual tokens for further reusing without requiring repeated compression for the same demonstration.
-We evaluate UniICL which only contains 17M trainable parameters on a scope of benchmarks involving linguistic acceptability, semantic classification, text summarization, and passage reranking, and UniICL achieves outstanding performances.
-Then, the in-domain experiments show that UniICL built upon two backbones (Vicuna and BlueLM) effectively substitutes the 12$\times$ demonstrations with soft prompts, easily scaling up the inner LLM from 4-shot to 64-shot.
+Existing work attempts to train the retrieval model and the generator in an end-to-end manner, which has shown better performance in in-domain datasets Wang et al.
+(2023c); Qiao et al.
+(2024).
+However, this approach still performs poorly in out-of-domain datasets.
+For instance, as shown in Fig.
+1(b), the retriever selects an example lexically similar to queries but has contrasting labels.
+Then, the LLM is misled and responds with a wrong answer.
+In light of challenges in ICL, we turn to leverage the inherent understanding ability of LLMs developed during pre-training.
+We accordingly propose a Unified ICL (UniICL) framework, which unifies demonstration compression, demonstration selection, and response generation.
+As shown in Fig.
+1(c), for lightweight training, in UniICL, both the compressor and generator are initialized from the same LLM and kept frozen.
+An adapter is introduced to align the compressor with the generator, and [M] is a learnable embedding called Memory Slot which is attached behind demonstrations for compression.
+Therefore, UniICL only contains 17M trainable parameters.
+The LLM compressor first compresses each demonstration from the training set and queries into Memory Tokens independently on top of Memory Slots.
+Then, UniICL selects $n$ most relevant demonstrations based on the similarity of Memory Tokens between queries and demonstrations.
+Finally, Memory Tokens of selected demonstrations are concatenated to formulate a global in-context sequence, together with queries fed into the generator for response generation.
+Due to independent compression, the compressor gets rid of the input window limitation of original LLMs as the number of demonstrations increases.
+In addition to improvements in window limitation, the tailored compression strategy further makes improvements to ICL efficiency.
+Specifically, UniICL caches Memory Tokens of different demonstrations to configure the Demonstration Bank (DB) for future reusing as shown in Fig.
+2.
+Therefore, repeated compression of the same demonstration is not necessary, which significantly boosts model efficiency in Fig.
+8.
+Extensive out-of-domain evaluation indicates UniICL achieves substantial improvements compared with other baselines.
 Our main contributions are as follows:
-  • To our knowledge, we are the first to propose an ICL framework that unifies compression, selection, and generation via a single frozen LLM.
-  • UniICL is a memory-friend framework that enables LLMs to perform large-shot ICL on consuming GPUs.
-  • UniICL proposes to configure Demonstration Bank to enhance ICL efficiency, avoiding repeated compression for the same demonstration.
+Figure 2: The workflow of Demonstration Bank.
+  • To our knowledge, we are the first to propose a unified ICL framework with 17M trainable parameters.
+  • UniICL proposes configuring the Demonstration Bank to avoid repeated compression for the same demonstration, which significantly boosts ICL efficiency.
+  • Different from the indiscriminate compression of previous studies, UniICL proposes a tailored compression strategy for ICL, achieving substantial improvements compared with other baselines.
 2 Related Work
@@
 To tackle the generalization problem, Mu et al.
-(2023) proposed to learn a Llama-7b to compress instruction to virtual tokens, but only compress instruction was not powerful enough since the demonstrations were much longer in practice.
-To compress the demonstrations, Chevalier et al.
+(2023) proposed to learn a Llama-7b to compress instructions to virtual tokens, but only compressing instructions was not powerful enough since the demonstrations were much longer in practice.
+To compress longer prompts, Chevalier et al.
 (2023) proposed AutoCompressor to recurrently generate compressed virtual tokens based on a fine-tuned Llama Zhang et al.
@@
 Apart from employing soft prompts, researchers also endeavored to shorten prompts by extracting informative tokens from the original ones Li (2023); Jiang et al.
-(2023), namely token pruning Kim et al.
+(2023), namely, token pruning Kim et al.
 (2022) or token merging Bolya et al.
@@
 (2023) to perform the alignment.
-However, UniICL naturally bypassed the distribution-aligned module, since the compressor and the target LLM were the same.
-Methods Additional Compressor Compression Tool # Trainable Parameters Train Size LLMLingua YES Pruning 7B 57k AutoCompressor NO Soft Prompt 7B UNKNOWN ICAE YES Soft Prompt 70M 240k UniICL NO Soft Prompt 17M 30k
-Table 1: Comparsion among recent compression methods and UniICL. Compression Tool represents the involved compression technique of different methods.
 3 Methodology
-We propose UniICL, a parameter-efficient ICL framework that unifies demonstration compression, demonstration selection, and response generation via a single LLM.
+Previous compression methods are not tailored for ICL, and they are either bound by serious inference latency or poor performance, as demonstrated in Appendix A.
+We propose UniICL, a unified ICL framework that unifies demonstration compression, demonstration selection, and response generation.
 As for the selection of the underlying LLM, previous work has proved that the Decoder-only model performs better than the Encoder-Decoder model in prompt compression Mu et al.
 (2023).
-We follow this conclusion and employ popular Vicuna-7b Zheng et al.
-(2023) and BlueLM-7B [^2] We mainly discuss UniICL built on Vicuna-7B, the experiments of BlueLM will be exhibited in Appendiex B.
-Team (2023) as the underlying backbone in UniICL.
-Figure 2: Differences of compression methods in formulating compressed virtual tokens in ICL.
-We present a comparison of training costs between UniICL and other recent compression methods in Table 1.
-Additionally, we illustrate differences in formulating virtual tokens for compression methods based on the soft prompt in Figure 2.
-To explain plainly, we ideally assume the compressor within three compression methods based on soft prompts has the window limitation of $L$, and has the same compression ratio, ignoring the length of soft prompts.
-In the 2-shot scenario, demonstrations $D_{1}$ and $D_{2}$ have a length of $L$.
-Considering the AutoCompressor, the concatenated demonstrations will be divided back into two segments, and the AutoCompressor compresses each segment step-by-step, bringing two times non-parallel compression.
-When it comes to ICAE, merely $D_{1}$ is accessible for the compressor and others will be read by no means.
-AutoCompressor shows advantages in the readable prompt length, but is short in efficiency, while ICAE has a constant compression complexity but struggles to approach relatively long inputs.
-Additionally, AutoCompressor makes the compression of later demonstrations be conditioned on the previous ones, which breaks demonstration independence in ICL.
-Combining the advantages of AutoCompressor and ICAE, UniICL compresses demonstrations independently and introduces virtual tokens concatenation to overcome long prompt challenges.
-In the $N$-shot settings, the number of practical compression steps can be calculated as $\lceil\frac{N}{k}\rceil$, where $k$ indicates that a single GPU is capable of compressing $k$ demonstrations in a batch.
-When the GPU capacity is sufficient, k equals $N$, which is the scenario of ICAE that compresses all segments in a time but UniICL drops nothing, while it degenerates to the AutoCompressor scenario that compresses segments step-by-step, when the GPU capacity is only sufficient to set $k=1$.
-Figure 3: Demonstration compression.
+We follow this conclusion and adopt Vicuna-7B Zheng et al.
+(2023) as the underlying backbone in UniICL.
+Figure 3: Demonstration compression. $k$ Memory Slots are attached behind each demonstration.
 3.1 Demonstration Compression
-UniICL introduces memory slots [M], a learnable embedding initialized from a rarely used embedding of the target LLM.
-UniICL activates the compression slots to absorb information from demonstrations in the forward propagation of frozen Vicuna, as illustrated in Figure 3.
-We first attach $k$ compression slots $=k\times\text{[M]}$M] to each demonstration $D_{i}$, formatting modified prompt fed to the Vicuna.
-Then, frozen Vicuna forwards the modified prompts and outputs the last hidden states $H_{i}=(h_{1},h_{2},...,h_{k})$ on top of the $k$ compression slots, dropping the others333$\oplus$ means token-level concatenation.:
-_$\_,H^{i}=\mathrm{forward}(D_{i}\oplus M)$ (1)
+[…]
+$H^{i}=f_{\theta}(D_{i}^{L_{i}\times d}\oplus M^{k\times d}),$ (1)
+where $L_{i}$ is the $i-$th demonstration length, $d$ is the embedding dimension and $\oplus$ means token-level concatenation.
 Due to the attention mechanism, $H^{i}$ is compelled to attend to the preceding actual tokens.
-Then, UniICL inserts a projection layer to convert $H^{i}$ into LLM-acceptable compressed virtual tokens $C^{i}=(c^{i}_{1},c^{i}_{2},...,c^{i}_{k})$:
-$c^{i}_{j}=W_{p}\cdot h^{i}_{j},$ (2)
+Then, UniICL applies a linear layer as the adapter for efficiency to convert $H^{i}$ to Memory Tokens $C^{i}=(c^{i}_{1},c^{i}_{2},...,c^{i}_{k})$, performing alignment between the compressor and the generator [^2] Linear layer is enough for UniICL as features have interacted with each other during compression.:
+$c^{i}_{j}=W_{p}^{d\times d}\cdot h^{i}_{j},$ (2)
 where $W_{p}$ is the parameters of the projection layer.
@@
 3.2 Demonstration Selection
-Except for substituting the origin demonstrations for generation, compressed virtual tokens $C^{i}$ are representative of demonstrations that can be again applied for demonstration selection, as illustrated in Figure 4.
-Specifically, given a query $Q$ and its candidate demonstrations $(D_{1},D_{2},...,D_{n})$, UniICL obtains their latent representation after average pooling:
-$\bar{C^{i}}_{Q/D}=\frac{1}{k}\sum_{j=1}^{k}c_{j}.$ (3)
-We define the $i$-th demonstration saliency score $S_{i}$ via the cosine similarity between Query and demonstrations:
-$S_{i}=\mathrm{cos}(\bar{C_{Q}},\bar{C}_{D}^{i}).$ (4)
-Demonstrations are then reranked according to their saliency scores.
-3.3 In-context Generation
-We employ the frozen Vicuna again to generate responses with the guiding of concatenated virtual demonstrations and queries, as illustrated in Figure 5.
-For $m$-shot in-context learning, we obtain $m$ virtual tokens after demonstration compression and selection, denoted as $C^{1}$ to $C^{m}$.
+Memory Tokens $C^{i}$ naturally summarize the demonstrations in latent space, and UniICL performs demonstration selection based on the similarity between queries and demonstrations as shown in Fig.
+4.
+Specifically, given a query $Q$ and its candidate demonstrations $(D_{1},D_{2},...,D_{n})$, UniICL obtains their representations used for selection by average pooling $C_{\{Q,D\}}$:
+$\bar{C^{i}}_{\{Q,D\}}=\frac{1}{k}\sum_{j=1}^{k}c_{j}.$ (3)
+We define the $i$-th demonstration saliency score $S_{i}$ as the cosine similarity between $\bar{C_{Q}}$ and $\bar{D_{i}}$:
+$S_{i}=\mathrm{cosine\_similarity}(\bar{C_{Q}},\bar{C}_{D}^{i}).$ (4)
+3.3 Generation
+We employ the frozen Vicuna again to generate responses with the guidance of concatenated Memory Tokens and queries, as illustrated in Fig.
+5.
+For $m$-shot in-context learning, we obtain $m$ spans of Memory Tokens after demonstration compression and selection, denoted as $C^{1}$ to $C^{m}$.
 Then, we horizontally concatenate them, keeping their relative position unmodified.
-Finally, the concatenated virtual tokens together with actual inference inputs are fed into Vicuna, performing auto-regressive generation as normal:
-$y_{i}=\mathrm{generate}(C^{1},...,C^{m};Q;y_{<i})$ (5)
-Except for the generative manner, virtual tokens are conveniently transferred to close-ended evaluation for understanding tasks through providing the candidate answers and measuring the label space PPL  [^4] https://huggingface.co/docs/transformers/perplexity, e.g.
-$(ppl^{+},ppl^{-}$) for sentiment classification:
-$y=\mathrm{argmin}(ppl^{+},ppl^{-}),$ (6)
-where answers with PPL closest to 1 are judged to be the current prediction.
-Figure 5: In-context generation.
+Finally, the concatenated Memory Tokens together with actual queries are fed into Vicuna, performing auto-regressive generation $g_{\theta}$ as normal:
+$y_{i}=g_{\theta}(C^{1},...,C^{m};Q;y_{<i}).$ (5)
+Except for the generative manner, Memory Tokens apply close-ended evaluation for understanding tasks as normal through measuring the perplexity of candidate choices  [^3] https://huggingface.co/docs/transformers/perplexity.
+Figure 5: In-context generation. The Memory Tokens from different demonstrations are concatenated horizontally at the input end of Vicuna.
 3.4 Training
-The trainable parameters in UniICL are merely 17M originating from the projection layer $W_{p}$ and the introduced compression slot [M].
-The projection layer is optimized with the language modeling objective of Vicuna to learn a base compression model.
+The trainable parameters in UniICL are merely 17M originating from the projection layer $W_{p}$ and the introduced Memory Slot [M].
+The linear layer is optimized with the language modeling objective $\mathcal{L}_{lm}$ of Vicuna to learn a base compression model.
 Then InfoNCE He et al.
@@
 $\mathcal{L}=\mathcal{L}_{lm}+\mathcal{L}_{ctr}.$ (6)
-Specifically, we slice the source input of each training instance into two parts and randomly compress one, denoted the compression part as $x_{c}$ and the unmodified as $x_{u}$.
-Afterward, we attach $M$ to $x_{c}$ and get virtual tokens $C$ on top of the compression slots, as described in Equ.
-1 and Equ.
+Specifically, we slice the source input of each training instance into two parts and randomly compress one.
+The compressed part is denoted as $x_{c}$ and the uncompressed part is denoted as $x_{u}$.
+Afterward, we attach the Memory Slot sequence $M$ behind $x_{c}$ and get Memory Tokens $C$ on top of the Memory Slots, as described in Eq.
+1 and Eq.
 2.
 Therefore, the language modeling loss $\mathcal{L}_{lm}$ is obtained as:
-$\mathcal{L}_{lm}=-\frac{1}{|y|}\sum_{t=0}logP(y_{t}|Q;C;y_{<t}),$ (8)
+$\mathcal{L}_{lm}=-\frac{1}{|y|}\sum_{t=0}logP(y_{t}|x_{u};C;y_{<t}),$ (7)
 where $y$ is the reference label of the current training instance.
 Additionally, to approach the large-shot settings without significant truncation, we introduce concatenation compression.
-When $x_{c}$ exceeds the window limitation for compression, UniICL further divides $x_{c}$ into acceptable ranges and compresses them independently to get local virtual tokens.
-Then, these virtual tokens of several segments will be concatenated to formulate global virtual tokens to replace $x_{c}$.
-Consequently, we utilize contrastive learning for selection augmentation and mine positives and negatives as illustrated in Figure 6.
-Specifically, given each training instance $Q$ and $n$ candidate demonstrations $(D_{1},D_{2},...,D_{n})$ from two non-crossing training subsets, we employ Vicuna to calculate the PPL concerning the golden label of $Q$, denoted as $ppl^{Q}$.
+When $x_{c}$ exceeds the window limitation for compression, UniICL further divides $x_{c}$ into acceptable ranges and compresses them independently to get local Memory Tokens.
+Then, these Memory Tokens from different segments will be concatenated to formulate global virtual tokens to replace $x_{c}$, applying Eq.
+7 to optimize models as well.
+We obtained a base compression model that has learned to compress and understand concatenated Memory Tokens after the first-phase training mentioned.
+Subsequently, we utilize contrastive learning for selection augmentation and mine positives and negatives as illustrated in Fig.
+6.
+Specifically, given each training instance $Q$ and $n$ candidate demonstrations $(D_{1},D_{2},...,D_{n})$ from two non-crossing training subsets, we employ Vicuna to calculate the PPL concerning the golden label of $Q$, denoted as $ppl^{Q}$ to find useful demonstrations for generation.
 Then, we provide the $i$-th demonstration and calculate PPL concerning the golden label of $Q$, denoted as $(ppl^{D}_{i},i\in[1,n])$.
@@
 $\widetilde{ppl}^{D}_{i}=ppl^{Q}-ppl^{D}_{i},i\in[1,n].$ (8)
-Hence, $C_{D}^{+}$ $(C_{D}^{-})$ is the representation of demonstrations $D^{+}$ ($D^{-}$) that furthest reduces (increases) $ppl^{Q}$ as processed in Equ.
-3, and the contrastive loss $\mathcal{L}_{ctr}$ can be formulated as:
-$\mathcal{L}_{ctr}=\frac{exp(cos(C_{Q},C_{D}^{+}))}{exp(cos(C_{Q},C_{D}^{+}))+exp(cos(C_{Q},C_{D}^{-}))}.$ (10)
+After finding demonstrations $D^{+}$ ($D^{-}$) that furthest reduces (increases) $ppl^{Q}$, we obtain their representation $C_{D}^{+}$ $(C_{D}^{-})$ as processed in Eq.
+3.
+The contrastive loss $\mathcal{L}_{ctr}$ can be formulated as:
+$\mathcal{L}_{ctr}=\frac{\exp(cos(C_{Q},C_{D}^{+}))}{\exp(cos(C_{Q},C_{D}^{+}))+\exp(cos(C_{Q},C_{D}^{-}))}.$ (9)
 In particular, if all relative PPL gains are less than 0, namely none of the candidate demonstrations help guide Vicuna to generate the golden label, we will apply the other set of candidates.
-Figure 6: Negatives mining pipeline.
+Figure 6: Contrastive examples mining pipeline. Finds demonstrations benefit/hinder the final generation according to the PPL.
 4 Experiment
 4.1 Baselines
-Naive Vicuna-7b serves as the fundamental baseline fed into actual demonstrations.
-AutoCompressor recurrently compresses demonstrations into virtual tokens.
-We employ their Llama2-7b version [^5] https://github.com/princeton-nlp/AutoCompressors..
+Unmodified Vicuna-7b serves as the fundamental baseline fed with actual demonstrations.
+AutoCompressor compresses prompts into 50 virtual tokens in different rounds recurrently.
+Previous compressed virtual tokens are put at the beginning of the current segment.
+Finally, virtual tokens of different compression rounds are concatenated for generation.
+We employ their Llama2-7b version for comparison.
 LLMLingua is a coarse-to-fine demonstration pruning method based on dropping uninformative words.
-We employ their released 7b version [^6] https://github.com/microsoft/LLMLingua., of which compressor is a fine-tuned Llama-2.
+We employ their released 7b version, of which the compressor is a fine-tuned Llama2.
 For a meaningful comparison, we replace target LLMs of LLMLingua (GPT-3.5-Turbo or Claude-v1.3) with the Vicuna-7b.
-ICAE [^7] https://github.com/getao/icae.
-compresses demonstrations into soft prompts via a LoRA-adapted Llama2-7b.
+ICAE compresses demonstrations into 128 virtual tokens via a LoRA-adapted Llama2-7b.
 Additionally, since selection augmentation is involved in the training of UniICL, we utilize the popular Sentence-BERT (S-BERT) Reimers and Gurevych (2019) as the dense retriever to construct an ICL pipeline for the above methods, serving as simple but effective selection-based baselines.
 4.2 Settings
-Considering the involved datasets and computation efficiency, we set the max allowed input length limit to 512 for both compression and generation.
-For a fair comparison, we set the allowed window of baselines to 512 and their compression ratio to the same as UniICL.
+Dataset # words (96,512] (512,1024] (1024,1536] XSum Narayan et al. (2018) - 10,000 4,697 CICERO Ghosal et al. (2022) 10,000 - - SUPER-NI Wang et al. (2022b) - 10,000 7,000 XSum (Ctr) 5,000
+Table 1: The composition training set of UniICL. (m,n] represents the range of the number of words in each instance. XSum (Ctr) is used for the second-phase training in Eq. 6.
+Dataset In-Domain # Test # Demonstrations MS MARCO-dev 6,980 - XSum 1,500 204,045/20 Arxiv 1,500 203,037/20 CoLA-dev 1,041 67,349/20 SST-2-dev 872 8,551/20 IMDb 1,500 25,000/20 MMLU 13,985 25,000/20
+Table 2: The details of the involved evaluation datasets.
+-dev represents employing the development set due to their test sets are inaccessible.
+# Demonstrations represent the number of demonstrations to be selected in high/low-resource ICL settings.
+We construct the training set by mixing up XSum, CICERO, and SUPER-NI according to their length as shown in Tab.
+1 and evaluate UniICL on extensive out-of-domain datasets as listed in Tab.
+2, with more details reported in Appendix H.
+Considering computation efficiency, we set the max allowed input length limit to 512 for both compression and generation for both training and inference.
+For a fair comparison, we set the allowed window of baselines to 512, and the compression ratio of default UniICL and baselines is set to 12, which is determined by the validation in Fig.
+7.
 We fix the learning rate to 8e-5 and use Adam as the optimizer, and the effective batch size is 32 (8 GPUs data parallelism and 4 steps gradient accumulation).
-Additionally, we conducted all experiments on 8*NVIDIA A5000 24G GPUs based on BFloat 16 data type, and we set the evaluated shot to 8 for understanding tasks and 5 for generative tasks for illustration because of marginal ICL gains and memory costs.
+We train 10 epochs and 2 epochs respectively for the first- and second-phase training.
+The best checkpoints are selected according to their performance on in-domain validation sets.
+Additionally, we conducted all experiments on 8*NVIDIA A5000 24G GPUs based on BFloat 16 data type, and we set the evaluated shots to 8 for understanding tasks and 5 for generative tasks for illustration, because of marginal ICL gains and memory costs.
 We apply S-BERT to pre-rank and output the top 10 similar candidates from training sets according to each inference input for all baselines.
-UniICL is employed to perform selection among them in practice due to computation efficiency for high-resource ICL.
-On the contrary, the low-resource ICL setting fixes the random candidate demonstrations to 20 for all inference inputs, performing pre-ranking and selecting as well.
+UniICL is employed to perform selection among them in practice due to computational efficiency for high-resource ICL.
+On the contrary, the low-resource ICL setting utilizes the randomly sampled 20 candidate demonstrations for all inference inputs, while UniICL performs selection as normal.
+To verify the universality, we further build UniICL on BlueLM-7B Team (2023) and Llama2-7B Touvron et al. (2023). Results of BlueLM and Llama2 will be reported in Appendix C and Appendix D.
 4.3 Results
 Model #-shots CoLA-dev SST-2-dev IMDb Arxiv XSum Acc.
-R-1 R-2 R-L R-1 R-2 R-L Vicuna 0-shot 56.2 91.7 92.6 34.3 9.1 27.4 19.9 5.0 13.4 1-shot 58.2 (57.4) 90.7 (90.8) 91.9 (91.0) 34.4 (33.2) 9.1 (8.5) 27.5 (26.7) 21.2 (20.4) 5.8 (5.2) 14.5 (13.9) 2-shot 62.1 (59.8) 92.1 (91.3) 91.7 (91.7) - - - - - - 5-shot 62.3 (61.9) 93.0 (91.9) 94.1 (92.5) - - - - - - AutoCompressor 1-shot 42.1 (40.9) 85.7 (84.2) 95.0 (95.1) 27.0 (26.4) 8.4 (8.2) 26.1 (25.8) 21.3 (20.3) 6.5 (6.3) 13.7 (13.7) 2-shot 58.8 (56.3) 88.0 (86.4) 95.0 (94.6) 27.1 (26.2) 8.6 (7.9) 26.4 (25.4) 21.9 (21.4) 6.6 (6.4) 14.5 (14.1) 5-shot 59.1 (58.8) 91.3 (89.1) 94.7 (94.8) 34.5 (33.7) 9.4 (9.1) 28.7 (27.9) 22.4 (21.7) 6.9 (6.7) 14.8 (14.3) LLMLingua 1-shot 55.5 (55.0) 89.7 (89.6) 91.0 (89.9) 33.3 (33.1) 8.9 (8.7) 27.4 (27.1) 20.5 (19.7) 5.4 (5.2) 14.5 (14.4) 2-shot 56.7 (55.7) 90.7 (90.2) 91.3 (91.0) 32.9 (32.0) 8.2 (8.1) 26.9 (25.9) 20.3 (20.0) 5.2 (5.1) 14.3 (14.1) 5-shot 57.2 (56.9) 90.6 (90.2) 90.9 (91.2) 30.1 (29.7) 7.9 (7.4) 25.3 (24.6) 19.7 (18.6) 4.9 (4.9) 14.1 (14.3) ICAE 1-shot 30.9 (30.9) 61.0 (60.1) 85.7 (83.3) 26.8 (24.6) 8.2 (7.1) 24.7 (22.9) 23.5 (21.9) 8.5 (7.8) 20.9 (20.3) 2-shot 30.9 (30.9) 49.0 (52.8) 85.9 (85.9) 27.2 (25.5) 8.4 (7.6) 25.9 (24.3) 24.4 (23.2) 8.9 (8.4) 21.3 (20.8) 5-shot 30.9 (30.9) 54.2 (51.0) 85.7 (85.9) 28.3 (26.9) 8.7 (7.7) 26.6 (25.8) 25.3 (24.9) 9.2 (8.8) 22.5 (21.6) UniICL 1-shot 58.7 (58.0) 92.9 (91.7) 94.3 (92.3) 35.5 (34.7) 10.5 (10.2) 28.7 (27.9) 27.7 (25.5) 10.2 (9.1) 21.2 (20.0) 2-shot 62.4 (61.0) 92.4 (91.6) 94.9 (93.3) 36.1 (35.2) 10.8 (10.4) 29.4 (28.2) 29.4 (26.8) 11.0 (9.8) 22.3 (20.9) 5-shot 62.6 (61.8) 93.1 (92.3) 94.5 (94.0) 35.8 (35.4) 10.6 (10.2) 29.5 (28.1) 30.7 (27.6) 11.3 (10.1) 22.8 (21.4) UniICL 1-shot 59.1 (58.7) 93.0 (91.9) 94.5 (91.6) 34.8 (34.7) 10.4 (10.3) 28.1 (27.8) 29.1 (26.2) 10.8 (9.4) 22.2 (20.7) 2-shot 62.6 (61.2) 94.0 (93.0) 94.9 (92.3) 34.6 (34.3) 10.6 (10.4) 28.5 (28.3) 30.3 (28.9) 11.3 (10.5) 22.9 (21.7) 5-shot 63.3 (61.5) 94.7 (92.8) 95.0 (93.8) 35.6 (35.3) 11.0 (10.8) 29.1 (27.7) 31.1 (30.0) 11.7 (11.2) 23.5 (22.3) 8-shot 63.8 (62.6) 94.7 (93.1) 95.0 (94.2) - - - - - - UniICL + $L_{ctr}$ 1-shot 59.3 (58.9) 93.2 (92.4) 95.1 (92.8) 35.6 (35.1) 10.7 (10.5) 28.9 (28.3) 30.0 (27.9) 11.3 (10.1) 22.8 (21.5) 2-shot 62.4 (62.0) 94.5 (92.8) 94.8 (93.4) 36.8 (35.3) 10.8 (10.6) 29.6 (28.9) 30.8 (29.2) 11.4 (10.7) 23.0 (21.9) 5-shot 64.3 (61.8) 94.7 (93.4) 96.1 (94.2) 37.1 (34.9) 11.3 (11.2) 30.0 (29.3) 32.5 (30.6) 12.3 (11.8) 24.7 (23.8) 8-shot 64.7 (63.3) 94.7 (94.1) 95.6 (95.0) - - - - - -
+R-1 R-2 R-L R-1 R-2 R-L Vicuna 0-shot 56.2 91.7 92.6 34.3 9.1 27.4 19.9 5.0 13.4 1-shot 58.2 (57.4) 90.7 (90.8) 91.9 (91.0) 34.8 (34.4) 9.3 (9.1) 27.9 (27.5) 21.5 (21.2) 5.9 (5.8) 14.7 (14.5) 2-shot 62.1 (59.8) 92.1 (91.3) 91.7 (91.7) - - - - - - 5-shot 62.3 (61.9) 93.0 (91.9) 94.1 (92.5) - - - - - - AutoCompressor 1-shot 42.1 (40.9) 85.7 (84.2) 95.0 (95.1) 27.0 (26.4) 8.4 (8.2) 26.1 (25.8) 21.3 (20.3) 6.5 (6.3) 13.7 (13.7) 2-shot 58.8 (56.3) 88.0 (86.4) 95.0 (94.6) 27.1 (26.2) 8.6 (7.9) 26.4 (25.4) 21.9 (21.4) 6.6 (6.4) 14.5 (14.1) 5-shot 59.1 (58.8) 91.3 (89.1) 94.7 (94.8) 34.5 (33.7) 9.4 (9.1) 28.7 (27.9) 22.4 (21.7) 6.9 (6.7) 14.8 (14.3) LLMLingua 1-shot 55.5 (55.0) 89.7 (89.6) 91.0 (89.9) 33.3 (33.1) 8.9 (8.7) 27.4 (27.1) 20.5 (19.7) 5.4 (5.2) 14.5 (14.4) 2-shot 56.7 (55.7) 90.7 (90.2) 91.3 (91.0) 32.9 (32.0) 8.2 (8.1) 26.9 (25.9) 20.3 (20.0) 5.2 (5.1) 14.3 (14.1) 5-shot 57.2 (56.9) 90.6 (90.2) 90.9 (91.2) 30.1 (29.7) 7.9 (7.4) 25.3 (24.6) 19.7 (18.6) 4.9 (4.9) 14.1 (14.3) ICAE 1-shot 30.9 (30.9) 61.0 (60.1) 85.7 (83.3) 26.8 (24.6) 8.2 (7.1) 24.7 (22.9) 23.5 (21.9) 8.5 (7.8) 20.9 (20.3) 2-shot 30.9 (30.9) 49.0 (52.8) 85.9 (85.9) 27.2 (25.5) 8.4 (7.6) 25.9 (24.3) 24.4 (23.2) 8.9 (8.4) 21.3 (20.8) 5-shot 30.9 (30.9) 54.2 (51.0) 85.7 (85.9) 28.3 (26.9) 8.7 (7.7) 26.6 (25.8) 25.3 (24.9) 9.2 (8.8) 22.5 (21.6) UniICL 1-shot 58.7 (58.0) 92.9 (91.7) 94.3 (92.3) 35.5 (34.7) 10.5 (10.2) 28.7 (27.9) 27.7 (25.5) 10.2 (9.1) 21.2 (20.0) 2-shot 62.4 (61.0) 92.4 (91.6) 94.9 (93.3) 36.1 (35.2) 10.8 (10.4) 29.4 (28.2) 29.4 (26.8) 11.0 (9.8) 22.3 (20.9) 5-shot 62.6 (61.8) 93.1 (92.3) 94.5 (94.0) 35.8 (35.4) 10.6 (10.2) 29.5 (28.1) 30.7 (27.6) 11.3 (10.1) 22.8 (21.4) UniICL 1-shot 59.1 (58.7) 93.0 (91.9) 94.5 (91.6) 34.8 (34.7) 10.4 (10.3) 28.1 (27.8) 29.1 (26.2) 10.8 (9.4) 22.2 (20.7) 2-shot 62.6 (61.2) 94.0 (93.0) 94.9 (92.3) 34.6 (34.3) 10.6 (10.4) 28.5 (28.3) 30.3 (28.9) 11.3 (10.5) 22.9 (21.7) 5-shot 63.3 (61.5) 94.7 (92.8) 95.0 (93.8) 35.6 (35.3) 11.0 (10.8) 29.1 (27.7) 31.1 (30.0) 11.7 (11.2) 23.5 (22.3) 8-shot 63.8 (62.6) 94.7 (93.1) 95.0 (94.2) - - - - - - UniICL + $L_{ctr}$ 1-shot 59.3 (58.9) 93.2 (92.4) 95.1 (92.8) 35.6 (35.1) 10.7 (10.5) 28.9 (28.3) 30.0 (27.9) 11.3 (10.1) 22.8 (21.5) 2-shot 62.4 (62.0) 94.5 (92.8) 94.8 (93.4) 36.8 (35.3) 10.8 (10.6) 29.6 (28.9) 30.8 (29.2) 11.4 (10.7) 23.0 (21.9) 5-shot 64.3 (61.8) 94.7 (93.4) 96.1 (94.2) 37.1 (34.9) 11.3 (11.2) 30.0 (29.3) 32.5 (30.6) 12.3 (11.8) 24.7 (23.8) 8-shot 64.7 (63.3) 94.7 (94.1) 95.6 (95.0) - - - - - -
 Table 3: The high- and low-ICL results on CoLA-dev, SST-2-dev, and IMDb.
-Results in () represent low-resource ICL.
+Results in (bracket) represent low-resource ICL.
 represents the demonstrations selected by UniICL, and the others are selected by S-BERT.
-+$L_{ctr}$ indicates the selection augmented UniICL (optimized with Equation 7).
++$L_{ctr}$ indicates the selection augmented UniICL (optimized with Eq.
+6).
 Bold (underline) represents the best performance on high- and low-resource ICL.
-We comprehensively evaluate the ICL performance of UniICL on the out-domain dataset CoLA, SST-2, and IMDb by close-ended evaluation and ARXIV by open-ended evaluation, as demonstrated in Table 2.
-Specifically, UniICL outperforms naive Vicuna-7b fed with actual candidate demonstrations, which indicates that virtual demonstrations are more efficient and informative for guiding the target LLM, and UniICL outperforms all the baselines by compressing the same demonstrations pre-ranked by S-BERT.
-Additionally, UniICL achieves further performance gains after selecting demonstrations via itself.
-Additionally, open-ended results indicate that virtual demonstrations still efficiently capture semantic information for ICL guiding, even though summarization demonstrations are much longer than understanding ones.
+R- indicates Rouge scores.
+All compression methods are evaluated with a compression ratio set to 12.
+We comprehensively evaluate the ICL performance of UniICL on the out-of-domain dataset CoLA, SST-2, and IMDb by close-ended evaluation and Arxiv by open-ended evaluation in Tab.
+3.
+The details of the involved evaluation datasets and metrics are reported in Tab.
+2 and Appendix H.
+Specifically, UniICL outperforms unmodified Vicuna-7b fed with actual candidate demonstrations, which indicates that Memory Tokens are more efficient and informative for guiding the target LLM.
+Meanwhile, UniICL outperforms all the baselines by compressing the same demonstrations pre-ranked by S-BERT.
+Additionally, UniICL achieves further performance gains after selecting demonstrations via itself (UniICL).
+The open-ended results highlight that Memory Tokens indeed capture semantic information for ICL generation, even though summarization demonstrations are much longer than understanding ones.
 Regarding Arxiv, the original ICL is not helpful enough due to its extremely over-length document, leaving little room for demonstrations.
-UniICL works as expected by compressing demonstrations and concatenating virtual demonstrations, and achieves +2.8 R-1 gains in the 5-shot setting that selects demonstrations via selection-augmented virtual tokens.
-Furthermore, The ablation experiments of +$\mathcal{L}_{ctr}$ show that UniICL is faced with performance degradation without $L_{ctr}$ and the performance gap becomes larger with the number of demonstrations increasing.
-The results of BlueLM are exhibited in Appendiex B.
+UniICL works as expected by compressing demonstrations into Memory Tokens and concatenating them, achieving +2.8 Rouge-1 gains in selection-augmented UniICL (+$\mathcal{L}_{ctr}$).
+Additionally, according to the results of +$\mathcal{L}_{ctr}$, we find that the gains brought by selection augmentation become larger as the number of demonstrations increases.
+We attribute this to the fact that UniICL selects more useful demonstrations for generation after the second-phase training.
+The results of BlueLM are exhibited in Appendix C.
+Except for understanding and generative tasks, we further evaluate UniICL on MMLU in Tab.
+4.
+UniICL achieves stable performance gains with more demonstrations introduced.
+Additionally, considering ICAE and AutoCompressor are soft-prompt-based compression methods built on Llama2, we also build UniICL on Llama2 for ablation in Appendix D.
+#-Shots S H SS O Avg. 0-shot 36.9 53.2 53.7 50.7 48.6 1-shot 38.6 55.3 54.6 52.4 50.2 2-shot 39.2 55.8 55.3 53.1 50.9 5-shot 40.1 55.6 55.3 53.8 51.2
+Table 4: Performance of UniICL on MMLU benchmark.
+We reported the Accuracy at the category level.
+S represents STEM, H represents Humanities, SS represents Social Science, O represents Other, and Avg indicates their average performance.
 Passage Ranking
-Method MS MARCO
-BM25 18.5
-Vicuna 28.9
-AutoCompressor 29.3
-ICAE 30.2
-UniICL 31.6
-Table 3: Results on MS MARCO.
+Method # TP MRR@10 BM25 - 18.5 Vicuna - 28.9 AutoCompressor - 29.3 ICAE - 30.2 UniICL - 31.6 SIMLM†‡ 110M 41.1 UniICL 17M 38.9
+Table 5: MRR@10 results on MS MARCO.
 Vicuna applies the last hidden states of [EOS] to represent sentences in latent space.
-Following the previous study, we report MRR@10.
-$\dagger$ means citing from Liang Wang et al.
-(2022).
-Since the virtual tokens naturally summarize semantic information of preceding sequences, we evaluate UniICL on the out-domain MS MARCO dataset in Table 3.
+Results citing from Liang Wang et al.
+(2022a) are denoted as , and methods supervised trained on MS MARCO are represented as .
+Bold indicates the best zero-shot performance and Underline is the best fine-tuned results.
+# TP indicates the number of trainable parameters.
+Since the virtual tokens naturally summarize semantic information of preceding sequences, we evaluate UniICL on the out-of-domain MS MARCO dataset in Tab.
+5.
 UniICL significantly outperforms the sparse retrieval method BM25 algorithm and other compression methods.
-Notably, we don’t compare UniICL with the popular retrieval models Reimers and Gurevych (2019); Wang et al.
-(2024) since most of them are fine-tuned on this dataset, which is unfair for comparison.
-Figure 7: The overall sensitivity analysis of compression ratio.
+Subsequently, we fine-tune the first-phase compression model of UniICL on the training set of MS MARCO.
+UniICL achieves comparable performance with SIMLM Wang et al.
+(2022a), which is specified in Information Retrieval (IR) and has more trainable parameters.
+Figure 7: The compression ratio sensitivity analysis of Llama2 , BlueLM, and Vicuna.
 5 Analysis
 5.1 Compression Ratio
+#-shots CoLA SST-2 IMDb Arxiv Acc.
+R-1 1-shot 58.5 (-0.8) 91.4 (-1.8) 92.6 (-2.5) 34.8 (-0.8) 2-shot 59.7 (-2.7) 92.1 (-2.4) 94.1 (-0.7) 35.7 (-1.1) 5-shot 62.4 (-1.9) 93.1 (-1.6) 94.8 (-1.3) 36.6 (-0.5)
+Table 6: Performance of UniICL on out-of-domain datasets, with a fixed compression ratio set to 12 during training.
 During training, the compression ratio is dynamically sampled from 2 to 16.
-We mix up 2,000 instances from the in-domain validation set, 1,000 for XSUM, and 1,000 for CICERO to select the compression ratio for UniICL in Figure 7, with the backbone of Vicuna and BlueLM respectively.
+We mix up 2,000 instances from the in-domain validation set, 1,000 for XSum, and 1,000 for CICERO to select the compression ratio for UniICL in Fig.
+7, with the backbone of Llama2, Vicuna, and BlueLM respectively.
 Specifically, UniICL compresses the latter cut-off part while keeping the former ones uncompressed.
@@
 However, when it comes to 16×16$\times$16 ×, an obvious drop occurs.
+In order to analyze this phenomenon more deeply, we provide a thorough analysis in Appendix G.
 Therefore, we set the compression ratio to 12 by default and apply this ratio to all experiments.
@@
 Memory explodes are represented as *, corresponding to the break of the line chart.
++Caching represents using DB.
+To explore whether it could yield additional performance gains compared with dynamic ratios, in Tab.
+6, we re-train UniICL with the compression ratio fixed to 12 (Results of more fixed ratios are reported in Appendix F.).
+Results indicate that UniICL trained with fixed compression ratios underperforms in out-of-domain datasets as it exhibits over-fitting in in-domain sets as shown in Tab.
+11.
+Furthermore, we analyze whether 12$\times$ is suitable for all out-of-domain datasets in Fig.
+9 in Appendix E.
+Results indicate that 12$\times$ outperforms other compression ratios in general across 4 out-of-domain datasets.
+It also points out that lower ratios still work comparable for short demonstrations and higher ratios are suitable for long demonstrations to some extent.
 Method GPUHours TFLOPs TMACs Vicuna 1.5 86,20 4,309 Vicuna-1k 1.9 31,664 15,832 UniICL 1.6 22,437 11,218
@@
 In UniICL, we incorporate an additional 17M trainable parameters into the 7b backbone, accounting for an approximate increase of 0.24%.
-We additionally evaluate the memory costs inference latency of UniICL and other compression methods in Figure 8.
+We evaluate the memory costs and inference latency of UniICL and other compression methods in Fig.
+8.
 With the help of the Demonstration Bank (DB), UniICL will eliminate the extra latency if the selected demonstrations have been compressed and cached (UniICL+Caching).
-Despite this, parallel computation facilitates the compressing process, resulting in minimal throughput degradation (UniICL and Baseline).
-What’s more, the naive 7B LLM occurs memory explosion for 8-shot settings and other compression methods perform up to 32-shot, while UniICL successfully scales up to 64-shots within 24GB CUDA allocation.
-Additionally, We demonstrate the inference computation and GPU hours in Table 4, by using 1,024 random legal tokens as inputs and forcing models to generate 128 tokens.
+Despite this, parallel computation facilitates the compression process, resulting in minimal throughput degradation (UniICL and Baseline).
+The unmodified 7B LLM causes a memory explosion for 8-shot settings, and other compression methods perform up to 32-shot, while UniICL successfully scales up to 64-shot within a 24GB CUDA allocation.
+Additionally, we demonstrate the inference computation and GPU hours in Tab.
+7, by using 1,024 random legal tokens as inputs and forcing models to generate 128 tokens.
 Notably, UniICL (without DB) compresses the former half, and the latter half is fed into the generator directly, while Vicuna and Vicuna-1k are distinguished in window limitations.
 Results indicate that minimal GPU hours increased due to the parallel computation of forward, although the extra compression of UniICL surges the computation.
-Additionally, Vicuna with a 1k window limitation surges both GPU hours and TFLOPs because long input brings significant computation and latency in generation.
-5.3 Training Deviation
-To quantify the performance gains brought by the learnable projection layer.
-We tune Vicuna and BlueLM with comparable parameters (17M) with LoRA in Table 5, setting the rank to 32.
-UniICL still outperforms LoRA-adapted LLMs with a 512 window limitation, indicating that the truncation indeed brings performance degradation.
+Additionally, Vicuna, with a 1k window limitation, surges both GPU hours and TFLOPs because long input brings significant computation and latency in generation.
 6 Conclusion
-This paper proposes UniICL, a parameter-efficient ICL framework that unifies demonstration selection, demonstration compression, and final response generation via a frozen LLM.
-Experimental results show that the generated virtual tokens substitute the 12$\times$ longer actual demonstrations with minimal time expenditure, scaling up the number of demonstrations from 4 to 64.
+This paper proposes UniICL, a parameter-efficient ICL framework that unifies demonstration selection, demonstration compression, and final response generation via a frozen LLM, an adapter, and a learnable embedding.
+Experimental results prove the advantages of UniICL in both efficiency and effectiveness.
+Due to 12$\times$ demonstration compression, UniICL scales up the number of demonstrations from 4 to 64 within a 24 GB VRAM allocation.
+Finally, to avoid repeated compression of the same demonstration, UniICL configures a Demonstration Bank (DB, which significantly boosts model efficiency.
 7 Limitations
 Our study, while proposing an efficient unified ICL framework for demonstration compression and selection, still has limitations.
-Firstly, UniICL is limited to the realm of naive ICL leaving other advanced LLM prompting methods, e.g.
-Retrieval Augment Generation (RAG) and Chain-of-Thought (CoT) unexplored.
-Limited to the hardware, we employ the underlying LLM at a scale of 7 billion parameters.
+Firstly, UniICL is limited to the realm of unmodified ICL, leaving other advanced LLM prompting methods, e.g., Retrieval Augment Generation (RAG) and Chain-of-Thought (CoT), unexplored.
+Limited to the hardware, we deploy the underlying LLM at a scale of 7 billion parameters.
 Larger-scale LLMs are welcome to enrich our findings in future studies.
-Appendix A In-Domain Evaluation
+8 Acknowledgement
+I would like to express my sincere gratitude to all the authors and reviewers for their valuable contributions to this research.
+This work was supported by the National Natural Science Foundation of China (NSFC 62106165) and the Project Funded by the Priority Academic Program Development of Jiangsu Higher Education Institutions, China.
+Appendix A Comparison with Existing Compression Methods
+Methods Additional Compression # Trainable Train Compressor Tool Parameters Size LLMLingua Jiang et al.
+(2023) YES Pruning 7B 57k AutoCompressor Wingate et al.
+(2022) NO Soft Prompt 7B UNKNOWN ICAE Ge et al.
+(2023) YES Soft Prompt 70M 240k UniICL NO Soft Prompt 17M 47k
+Table 8: Comparison among recent compression methods and UniICL.
+Compression Tool represents the involved compression technique of different methods.
+Train Size represents the size of the training datasets.
+We present a comparison of training costs between UniICL and other recent compression methods in Tab. 8.
+Appendix B In-Domain Evaluation
 Backbone Method XSum CICERO R-1 R-2 R-L R-1 R-2 R-L Vicuna-7b Vicuna 19.9 5.0 13.5 17.3 3.3 14.3    +LoRA 25.4 7.5 17.3 28.1 10.5 25.6 Vicuna-1k 27.3 8.7 19.7 30.5 11.3 27.4    +LoRA 31.2 11.0 23.1 34.1 13.5 30.2 UniICL 30.0 10.2 22.3 32.6 12.2 28.8 BlueLM-7b BlueLM 15.0 3.6 10.4 17.6 3.1 15.0    +LoRA 23.1 7.6 17.4 21.9 7.8 19.8 BlueLM-1k 28.1 9.9 22.8 25.1 9.2 23.1    +LoRA 30.8 10.5 24.6 31.2 10.8 27.4 UniICL 30.4 10.2 23.7 29.2 10.0 26.6
-Table 5: The in-domain results and ablation studies on XSUM and CICERO. 1k represents the extending 1k window limitation, while others have a limitation of 512.
-We conduct the zero-shot in-domain generation evaluation on the entire test set of XSUM and CICERO in Table 5 by compressing the latter half to virtual tokens and keeping the former unmodified.
-UniICL significantly outperforms the baselines, indicating the compressed virtual tokens can provide the original truncated information by recovering the cut-off parts after supervised fine-tuning.
+Table 9: The in-domain results and ablation studies on XSum and CICERO. 1k represents the extended 1k window limitation, while others have a limitation of 512.
+We conduct the zero-shot in-domain generation evaluation on the entire test set of XSum and CICERO in Tab.
+9 by compressing the latter half to virtual tokens and keeping the former unmodified.
+UniICL significantly outperforms the baselines, indicating that the compressed virtual tokens can provide the original truncated information by recovering the cut-off parts after supervised fine-tuning.
 Although extending the window to 1k, Vicuna and BlueLM still underperform UniICL, indicating that compressed virtual tokens filter noise information to some extent.
-Appendix B Results on BlueLM
-We extra conduct experiments on BlueLM Team (2023) to verify the generality of UniICL. We demonstrate the result of understanding tasks in Table 6, of the generative tasks in Table 7.
+Additionally, to quantify the performance gains brought by the learnable projection layer.
+We tune Vicuna and BlueLM with comparable parameters (17M) with LoRA, setting the rank to 32 in Tab.
+9.
+UniICL still outperforms LoRA-adapted LLMs with a 512 window limitation, indicating that the truncation indeed brings performance degradation.
+Appendix C Results on BlueLM
+We also conduct experiments on BlueLM Team (2023) to verify the generality of UniICL. We demonstrate the result of understanding tasks in Tab. 10, of the generative tasks in Tab. 11.
 Model #-shots CoLA-dev SST-2-dev IMDb Acc.
-BlueLM 0-shot 71.6 81.2 48.8 1-shot 69.6 82.6 64.8 2-shot 70.0 87.0 65.6 5-shot 70.5 88.6 68.7 UniICL 1-shot 69.6 81.2 65.4 2-shot 68.7 82.6 67.0 5-shot 71.7 87.0 70.4 UniICL 1-shot 69.8 80.0 62.0 2-shot 70.1 80.8 67.0 5-shot 71.8 85.6 69.6 8-shot 72.3 87.4 69.4 UniICL + $L_{ctr}$ 1-shot 70.1 80 69.6 2-shot 70.3 87.2 70.6 5-shot 71.1 89.2 71.0 8-shot 72.5 90.4 76.8
+BlueLM 0-shot 71.6 81.2 48.8 1-shot 69.6 82.6 64.8 2-shot 70.0 87.0 65.6 5-shot 70.5 88.6 68.7 UniICL 1-shot 69.6 81.2 65.4 2-shot 70.1 82.6 67.0 5-shot 71.8 87.0 70.4 UniICL 1-shot 70.1 80.0 62.0 2-shot 70.3 80.8 67.0 5-shot 71.1 85.6 69.6 8-shot 71.6 87.4 69.4 UniICL + $L_{ctr}$ 1-shot 68.9 80.0 69.6 2-shot 70.6 87.2 70.6 5-shot 71.5 89.2 71.0 8-shot 72.4 90.4 76.8
 Table 10: The ICL results of understanding tasks with the backbone of BlueLM.
-Method #-shots XSUM ARXIV R-1 R-2 R-L R-1 R-2 R-L BlueLM 0-shot 15.0 3.6 10.4 30.9 7.7 24.7 1-shot 19.1 4.8 12.1 23.0 3.6 19.0 UniICL 1-shot 24.0 6.9 18.0 31.4 7.7 25.2 2-shot 25.0 7.3 18.8 30.8 7.3 24.8 5-shot 25.0 7.3 18.8 31.9 7.8 26.0 UniICL 1-shot 25.2 7.4 18.9 31.6 7.9 25.4 2-shot 25.4 7.6 19.1 31.9 8.0 25.6 5-shot 26.5 7.9 20.3 32.1 8.0 25.5
+Method #-shots XSum Arxiv R-1 R-2 R-L R-1 R-2 R-L BlueLM 0-shot 15.0 3.6 10.4 30.9 7.7 24.7 1-shot 19.1 4.8 12.1 23.0 3.6 19.0 UniICL 1-shot 24.0 6.9 18.0 31.4 7.7 25.2 2-shot 25.0 7.3 18.8 30.8 7.3 24.8 5-shot 25.3 7.4 19.1 31.9 7.8 26.0 UniICL 1-shot 24.7 7.2 18.5 31.0 7.5 24.9 2-shot 25.1 7.4 19.0 31.2 7.7 25.1 5-shot 26.3 7.6 20.0 31.5 7.9 25.3 UniICL + $L_{ctr}$ 1-shot 25.2 7.4 18.9 31.6 7.9 25.4 2-shot 25.4 7.6 19.1 31.9 8.0 25.6 5-shot 26.5 7.9 20.3 32.1 8.0 25.5
 Table 11: The ICL results of generative tasks with the backbone of BlueLM.
-Appendix C Datasets & Metrics
-C.1 Datasets
-Dataset # words (96,512] (512,1024] (1024,1536] XSUM - 10,000 4,697 CICERO 10,000 - - SUPER-NI - 10,000 7,000 XSUM (Ctr) 5,000
-Table 8: The composition training set of UniICL. (m,n] represents the range of the number of words in each instance. XSUM (Ctr) is used for the second phase training in Equation 7.
-Dataset Task In-Domain # Test # Demonstrations MS MARCO-dev Passage Ranking 6,980 - XSUM Text Summarization 1,500 204,045/20 ARXIV Text Summarization 1,500 203,037/20 CoLA-dev Lingustic Acceptability 1,041 67,349/20 SST2-dev Sentiment Classification 872 8,551/20 IMDb Sentiment Classification 1,500 25,000/20
-Table 9: The details of involved evaluation datasets.
--dev represents employing development set due to their test sets are inaccessible.
-# Demonstrations represent the number of demonstrations to be selected in high/low-resource ICL settings.
-We mix up two public datasets for compression and selection augmentation training, described in Table 8.
-Additionally, UniICL achieves outstanding performance on out-domain evaluation, involving text summarization Narayan et al.
+Appendix D Supplementary Ablation on Llama2
+AutoCompressor Wingate et al.
+(2022) and ICAE Ge et al.
+(2023) are built on Llama2-7B Touvron et al.
+(2023), which are soft-prompt-based methods similar to UniICL.
+Therefore, we evaluate UniICL with Llama2 as the backbone.
+As shown in Tab 12 and Tab.
+13, UniICL achieves substantial improvements compared with unmodified Llama2 and outperforms ICAE and AutoCompressor demonstrated in Tab.
+3.
+Model #-shots CoLA-dev SST-2-dev IMDb Acc.
+Llama2 0-shot 73.4 93.0 85.3 1-shot 74.8 94.0 85.5 2-shot 75.6 94.9 87.8 5-shot 84.3 97.2 92.7 UniICL 1-shot 74.9 94.1 94.6 2-shot 75.9 95.1 96.1 5-shot 85.4 95.7 96.5
+Table 12: The ICL results of understanding tasks with the backbone of Llama2.
+Method #-shots XSum Arxiv R-1 R-2 R-L R-1 R-2 R-L Llama2 0-shot 27.4 7.6 20.1 32.9 8.9 29.2 1-shot 27.7 7.9 20.3 30.1 8.0 28.4 UniICL 1-shot 27.8 8.0 20.5 33.3 9.2 29.7 2-shot 28.4 8.6 21.3 34.0 9.4 30.3 5-shot 29.3 9.1 22.0 34.5 9.7 30.8
+Table 13: The ICL results of generative tasks with the backbone of LLama2.
+Appendix E Compression Ratio Selection on Different Tasks
+Figure 9: Winrate of different compression ratios on out-of-domain evaluation in 1-shot settings.
+We illustrate suitable ratio selection across four out-of-domain datasets in Fig.
+9.
+For tasks with relatively short inputs, such as CoLA and SST2, UniICL tends to perform better with a compression ratio set to 4.
+While in IMDb and Arxiv, which are longer, UniICL performs better with higher compression ratios.
+UniICL with a 12$\times$ compression ratio substantially outperforms other settings on four datasets.
+Figure 10: Winrate of UniICL with a fixed number of Memory Tokens.
+Additionally, we are curious about whether it is necessary to introduce more demonstrations with a higher compression ratio.
+In Fig.
+10, we find that the performance of compressing 2 demonstrations with a 12×12$\times$12 × ratio is stable and outperforms other settings across 3/4 datasets.
+6$\times$ compression ratio with 1 demonstration compressed performs worst in general.
+When compressing 4 demonstrations with a 24×24$\times$24 × ratio, its performance is comparable, and it slightly outperforms the 12×12$\times$12 × ratio in SST2.
+Appendix F Fixed Compression Ratio Training
+Figure 11: The relative performance on in-domain and out-of-domain datasets, with UniICL trained with a fixed ratio. Out-of-domain evaluation applies 1-shot settings.
+To verify the effectiveness of the dynamic sampled compression ratio of UniICL, we train models with more extensive fixed compression ratios and perform out-of-domain evaluation with the same ratio in Fig.
+11.
+Results indicate that fixed compression ratios work better than dynamic sampled ratios in in-domain evaluation, but underperform in out-of-domain evaluation.
+We attribute this to the fixed compression ratio makes models exhibit over-fitting during training, and demonstration compression degrades to Prefix Tuning.
+Appendix G Visualization of Memory Tokens
+Figure 12: Cosine similarity between Memory Tokens (vertical axis) and original embeddings (horizon axis).
+Figure 13: Attention scores on Memory Tokens in the first step generation.
+The vertical axis describes the 32 LLM layer, and the horizon axis indicates the number of Memory Tokens across different compression ratios.
+Above each figure, %/MT represents the average proportion of the attention score occupied by memory tokens in each LLM layer.
+To explore how Memory Tokens work within UniICL across different compression ratios, we visualize the cosine similarity between Memory Tokens and original embeddings in Fig.
+12 and attention scores of the first generation step in Fig.
+13.
+Intuitively, the 4$\times$ compression ratio should retain more information due to more Memory Tokens.
+However, as shown in Fig.
+12(a), the cosine similarity is relatively sparser than the 4$\times$ compression ratio illustrated in Fig.
+12(b).
+This tendency is aligned with the first step attention scores in Fig.
+13(a).
+According to merely 0.3% average attention occupied in a generation, we can conclude that more Memory Tokens fail to provide models with more information.
+We attribute this phenomenon to the given semantic information being distributed over all Memory Tokens as models attend to each Memory Token equally in Fig.
+13(a).
+Fewer Memory Tokens are enough to concentrate this information, represented as relatively concentrated similarity distribution in Fig.
+12(b) and higher attention scores in Fig.
+13(b), both of which indicate denser information retained.
+When the compression ratio becomes higher, such as 16 or 32, Memory Tokens become fewer and therefore sparse information retrained as shown in Fig.
+12(c), Fig.
+12(c), Fig.
+13(c), and Fig.
+13(c).
+This also provides an explanation for the slow performance degradation with ratios varying from 4 to 12 and drops sharply at 16 in Fig.
+7.
+Appendix H Datasets & Metrics
+Datasets
+We mix up three public datasets for compression and selection augmentation training, described in Tab.
+1.
+The training set includes an instruction dataset, SUPER-NI, which we used to make UniICL respond to various instructions.
+Notably, we don’t perform an in-domain evaluation on SUPER-NI as it only contains a training set.
+After training, we extensively evaluate UniICL on out-of-domain evaluation, involving text summarization Narayan et al.
 (2018), passage ranking Nguyen et al.
@@
 (2011); Socher et al.
-(2013), and linguistic acceptability Warstadt et al.
-(2018), more details referring to Table 9.
+(2013), linguistic acceptability Warstadt et al.
+(2018), and a popular reasoning benchmark Hendrycks et al.
+(2020), more details referring to Tab.
+2.
+MS MARCO is popularly used in Information Retrieval (IR), we use this dataset to evaluate the ability of UniICL to capture document-level information.
+Specifically, MMLU (Massive Multitask Language Understanding) is a comprehensive evaluation benchmark including 57 subjects from STEM, Humanities, Social Sciences, and Other fields.
+We use this benchmark to evaluate the reasoning ability of UniICL.
 UniICL selects demonstrations from its training set in high-resource ICL, and we fixed the number of candidate demonstrations to 20 for low-resource ICL evaluation.
-C.2 Evaluation Metrics
-ROUGE Lin (2004) is a widely adopted metric in many generative tasks that evaluate how similar the generated hypothesis is to the golden label.
+Evaluation Metrics
+ROUGE Lin (2004) is a widely adopted metric in many generative tasks that evaluates how similar the generated hypothesis is to the golden label.
 Therefore, ROUGE is used in our experiments to evaluate the quality responses generated conditioned on compressed virtual tokens, and we report the F-1 scores of ROUGE-1, ROUGE-2, and ROUGE-L (abbreviated R-1, R-2, R-L in the following), and we employed the files2rouge  [^4] https://github.com/pltrdy/files2rouge.
-library in practice. Following the previous works, we report the accuracy of close-ended evaluation and MRR@10 for passage ranking.
+library in practice.
+Following the previous works, we report the accuracy of close-ended evaluation and MRR@10 for passage ranking.